### v1

# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with $L_{2}$-mean Hörmander conditions

Abstract.
We construct a slightly new noncommutative Calderón-Zygmund decomposition by further splitting the bad function. Using this tool, we prove the weak type (1,1) boundedness of noncommutative Calderón-Zygmund operators under $L_{2}$-mean Hölder conditions, which improves the previous result.
Key words and phrases:
Noncommutative Calderón-Zygmund decomposition, Noncommutative Calderón-Zygmund operators
This work is supported by National Natural Science Foundation of China (No. 12322107, No. 12271124, and No. W2441002) and Heilongjiang Provincial Natural Science Foundation of China (No. YQ2022A005).
1. Introduction
Calderón-Zygmund operators, a well-known class of singular integral operators, were introduced by Calderón and Zygmund [Calderon1952On] in their seminal 1952 work. The kernel of such an operator satisfies specific conditions, including the size and regularity conditions, the latter of which was refined by Hörmander [Hormander1960Estimates] in 1960.
The standard nonconvolution type Calderón-Zygmund operator is defined as
$Tf(x)=\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)f(y){\rm d}y,\quad x\notin\operatorname{supp}f,$
where $K:\operatorname{\mathbb{R}}^{d}\times\operatorname{\mathbb{R}}^{d}\setminus\left\{(x,x):x\in\operatorname{\mathbb{R}}^{d}\right\}\rightarrow\operatorname{\mathbb{C}}$ is a standard nonconvolution type Calderón-Zygmund kernel, which is a binary function satisfying
  • Size condition: there exists $C_{1}>0$ such that
    (1.1) $|K(x,y)|\leq\frac{C_{1}}{|x-y|^{d}};$
  • Lipschitz regularity condition: there exists $\gamma\in[0,1]$ and $C_{2}>0$ such that
    (1.2) $|K(x,y)-K(x,z)|+|K(y,x)-K(z,x)|\leq\frac{C_{2}|y-z|^{\gamma}}{|x-y|^{d+\gamma}},\quad\text{if}\ |x-y|\geq 2|y-z|,$
where $x$, $y$, $z\in\operatorname{\mathbb{R}}^{d}$. The Lipschitz regularity condition (1.2) can be relaxed to certain variants of smoothness conditions (see [Grafakos2019Alimited, Suzuki2021TheCalderon, Watson1990Weighted]). In the following, all Calderón-Zygmund kernels under consideration are of nonconvolution type.
Inspired by operator algebras, harmonic analysis, noncommutative geometry, and quantum probability (see [Cadilhac2018Weak, Chen2013Harmonic, Junge2014Smooth, Junge2018Noncommutative, Mei2007Operator, Mei2009Pseudo-localization, Mei2017Free, Mei2022Mikhlin, Parcet2009Pseudo-localization]), noncommutative harmonic analysis has become an exciting field in recent years. However, generalizing the Calderón-Zygmund decomposition, a key technique in harmonic analysis first introduced in [Calderon1952On], to the noncommutative setting is challenging.
In 2009, Parcet [Parcet2009Pseudo-localization] rigorously constructed a kind of noncommutative Calderón-Zygmund decomposition by using Cuculescu’s maximal weak type $(1,1)$ estimates for noncommutative martingales [Cuculescu1971Martingales]. With this tool, Parcet established the weak type $(1,1)$ estimates for operator-valued Calderón-Zygmund singular integrals with the kernel satisfying the Lipschitz regularity condition (1.2) (Cadilhac later gave a shorter proof in [Cadilhac2018Weak]). Notably, the Hörmander condition
(1.3) $\sup_{y\in\operatorname{\mathbb{R}}^{d},|v|>0}\int_{|x-y|\geq 2|v|}|K(x,y+v)-K(x,y)|{\rm d}x<\infty,$
which is weaker than (1.2), is already sufficient for weak type $(1,1)$ boundedness in classical Calderón-Zygmund theory. It is therefore natural to ask whether the noncommutative weak type $(1,1)$ boundedness still holds under the Hörmander condition (1.3)? This question remains open.
A step forward was made in 2022 by Cadilhac et al. [Cadilhac2022Spectral]. They developed a more effective noncommutative Calderón-Zygmund decomposition without the off-diagonal term of the good function, which allowed them to prove the noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators if the kernel $K$ satisfies
(1.4) $\sum_{m\geq 1}\sup_{{\begin{subarray}{c}y\in Q\\ Q\ \text{dyadic}\end{subarray}}}\Big(2^{md}\ell(Q)^{d}\int_{2^{m}\ell(Q)\leq|x-c(Q)|\leq 2^{m+1}\ell(Q)}|K(x,y)-K(x,c(Q))|^{2}{\rm d}x\Big)^{\frac{1}{2}}<\infty,$
where $\ell(Q)$ and $c(Q)$ stand for the length and the center of the dyadic cube $Q$, respectively. More generally, for $1\leq q\leq\infty$, we say that a kernel $K$ satisfies the $L_{q}$-integral regularity condition if
(1.5) $\sum_{m\geq 1}\delta_{q}(m)<\infty,$
where $\delta_{q}(m)$ is defined as
$\sup_{{\begin{subarray}{c}y\in\operatorname{\mathbb{R}}^{d},\ R>0\\ |v|<R\end{subarray}}}$ $\Big((2^{m}R)^{d(q-1)}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x\Big)^{\frac{1}{q}},$
for $m\geq 1$. Roughly speaking, the Hörmander condition (1.3) corresponds to the $L_{1}$-integral regularity condition i.e. (1.5) with $q=1$, while (1.4) corresponds to the $L_{2}$-integral regularity condition, i.e., (1.5) with $q=2$. The fact $\delta_{1}(m)\lesssim_{d}\delta_{2}(m)$ shows that the $L_{2}$-integral regularity condition is stronger than the Hörmander condition (1.3), but weaker than the Lipschitz regularity condition (1.2). By applying the noncommutative Calderón-Zygmund decomposition in [Cadilhac2022Spectral], Hong, the first author, and Xu [hong2023maximal] established the weak type (1,1) boundedness for the noncommutative maximal truncated Calderón-Zygmund operators under the $L_{2}$-integral regularity condition. We refer the reader to [Cadilhac2022Spectral, Lai2024Noncommutative, Mei2009Pseudo-localization] for more results on weak (1,1) boundedness theory for noncommutative Calderón-Zygmund operators.
In this paper, we study a kernel smoothness condition that is slightly weaker than the $L_{q}$-integral regularity condition (1.5): the $L_{q}$-mean Hörmander condition ($1\leq q\leq\infty$), defined by
(1.6) $\sum_{m\geq 1}H_{q}(m)<\infty,$
where $H_{q}(m)$ is defined as
$\sup_{{\begin{subarray}y\in\operatorname{\mathbb{R}}^{d},\\ R>0\end{subarray}}}$ $\Big(\frac{(2^{m}R)^{d(q-1)}}{|B_{R}|}\int_{|v|\leq R}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x{\rm d}v\Big)^{\frac{1}{q}},$
for $m\geq 1$. Here, $B_{R}$ denotes the ball centered at the origin with radius $R>0$, and $|B_{R}|$ denotes its volume. It is not hard to see that $H_{q}(m)\lesssim_{d}\delta_{q}(m)$, meaning the $L_{q}$-mean Hörmander condition (1.6) is indeed weaker than the $L_{q}$-integral regularity condition (1.5). In what follows, we focus on the case $q=2$, i.e., the $L_{2}$-mean Hörmander condition:
(1.7) $\sum_{m\geq 1}H_{2}(m)<\infty,$
where $H_{2}(m)$ is defined as in (1.6) with $q=2$.
For two smoothness conditions $A$ and $B$, we write $A\prec B$ to mean that $A$ is weaker than $B$. Thus, we have the following conclusion:
[…]
This naturally leads to a question: can the $L_{2}$-mean Hörmander condition (1.7) still guarantee the noncommutative weak type $(1,1)$ boundedness of the maximal Calderón-Zygmund operators? This paper gives a positive answer to this question.
[…]
(1.8) $Tf(x)=\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)f(y){\rm d}y,\quad x\notin\overrightarrow{\operatorname{supp}}\ f,$
where $f\in L_{1}(\operatorname{\mathcal{N}})\bigcap L_{\infty}(\operatorname{\mathcal{N}})$ is compactly supported and measurable. Here, $\overrightarrow{\text{supp}}\ f$ denotes the support of $f$ as an operator-valued function in $\operatorname{\mathbb{R}}^{d}$ rather than the support projection as an element in a von Neumann algebra. For any $\epsilon>0$, the associated truncated singular integral $T_{\epsilon}f$ is defined as
(1.9) $T_{\epsilon}f(x)=\int_{|x-y|>\epsilon}K(x,y)f(y){\rm d}y.$
The main result is stated as follows.
Theorem 1.1.
Suppose $T$ is a Calderón-Zygmund operator defined as in (1.8) associated with a kernel satisfying (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}$ be defined as in (1.9). Suppose $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $1<p_{0}<\infty$, that is, for any $f\in L_{p_{0}}(\operatorname{\mathcal{N}})$,
$\left\|(T_{\epsilon}f)_{\epsilon>0}\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}.$
Then, $(T_{\epsilon})_{\epsilon>0}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
$\left\|(T_{\epsilon}f)_{\epsilon>0}\right\|_{\Lambda_{1,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We refer to see Appendix A for the definitions of the noncommutative maximal norm $\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{N}};\ell_{\infty})}$ and the noncommutative weak maximal norm $\left\|\cdot\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}$.
Our main strategy in this paper is to develop a refined noncommutative Calderón-Zygmund decomposition, adapted to kernels satisfying the $L_{2}$-mean Hörmander condition (1.7). The previous methods are insufficient in this context. On one hand, Parcet’s decomposition [Parcet2009Pseudo-localization] requires the stronger Lipschitz regularity condition (1.2) and may not be directly applied to the noncommutative maximal operator. On the other hand, while the decomposition by Cadilhac et al. [Cadilhac2022Spectral] can deal with the $L_{2}$-integral regularity condition (1.5) with $q=2$, it does not cover our target. To address this gap, we introduce a slightly new noncommutative Calderón-Zygmund decomposition achieved by further splitting of the bad function into a convolution part and a remainder part.
This paper is organized as follows. We begin by presenting our key tool: the refined noncommutative Calderón-Zygmund decomposition (Theorem 2.4) in Section 2. Next, Section 3 starts the proof of the main result (Theorem 1.1), reducing it to proving the weak type (1,1) estimates for the lacunary sequences with real kernels (Theorem 3.3). With this reduction established, Section 4 is devoted to proving Theorem 3.3. Finally, some related background concepts are included in Appendix A.
Now, we conclude this section by listing the notation required for our later analysis.
Notation
  • $\operatorname{\mathcal{Q}}_{n}$: the set of all dyadic cubes of side length $2^{-n}$, $n\in\operatorname{\mathbb{Z}}$.
  • $Q_{x,n}$: the unique cube in $\operatorname{\mathcal{Q}}_{n}$ containing $x\in\operatorname{\mathbb{R}}^{d}$.
  • $\operatorname{\mathcal{Q}}$: the set of all standard dyadic cubes in $\operatorname{\mathbb{R}}^{d}$, that is, $\operatorname{\mathcal{Q}}=\bigcup\limits_{n}\operatorname{\mathcal{Q}}_{n}$.
  • $c(Q)$: the center of the cube $Q$.
  • lQlQ: the concentric cube sharing the center of $Q$ such that its length is $l$ times the length of $Q$.
  • $B_{r}(x)$: the ball with $x\in\operatorname{\mathbb{R}}^{d}$ as its center and $r>0$ as its radius, and the center $x$ will be omitted when it is the origin.
  • $|E|$: the volume of set $E\subset\operatorname{\mathbb{R}}^{d}$.
  • $[k]$: the integer part of $k$.
2. Noncommutative Calderón-Zygmund decomposition
The Calderón-Zygmund decomposition is a fundamental tool for studying the boundedness of singular integral operators. When transferring to the von Neumann algebraic setting, its noncommutative counterpart is primarily based on Cuculescu’s martingale projections [Cuculescu1971Martingales]. In this section, we first review Cuculescu’s construction, and then use it to develop our refined version of the noncommutative Calderón-Zygmund decomposition. Background on noncommutative $L_{p}$ spaces and maximal norms can be found in Appendix A.
2.1. Cuculescu’s construction
Let $\operatorname{\mathcal{N}}=L_{\infty}(\operatorname{\mathbb{R}}^{d})\bar{\otimes}\operatorname{\mathcal{M}}$ be the von Neumann algebra tensor product, equipped with the tensor trace $\phi=\int_{\operatorname{\mathbb{R}}^{d}}\otimes\tau$. For $n\in\mathbb{Z}$, let $\sigma_{n}$ be the $\sigma$-algebra generated by $\operatorname{\mathcal{Q}}_{n}$, and let $\operatorname{\mathcal{N}}_{n}=L_{\infty}(\operatorname{\mathbb{R}}^{d},\sigma_{n},dx)\bar{\otimes}\operatorname{\mathcal{M}}$, which is a von Neumann subalgebra of $\operatorname{\mathcal{N}}$. Then, $(\operatorname{\mathcal{N}}_{n})_{n\in\mathbb{Z}}$ forms a sequence of increasing von Neumann subalgebras of $\operatorname{\mathcal{N}}$ whose union is weak-* dense in $\operatorname{\mathcal{N}}$.
Let $E_{n}$ be a conditional expectation from $\operatorname{\mathcal{N}}$ onto $\operatorname{\mathcal{N}}_{n}$. A sequence $(a_{n})_{n\in\operatorname{\mathbb{Z}}}\subset\operatorname{\mathcal{N}}$ is said to be a martingale if $E_{n}(a_{n+1})=a_{n}$. For $1\leq q\leq\infty$, if in addition $(a_{n})_{n\in\operatorname{\mathbb{Z}}}\subset L_{q}(\operatorname{\mathcal{N}})$, then $(a_{n})_{n\in\operatorname{\mathbb{Z}}}$ is called an $L_{q}$-martingale. In this paper, we define the conditional expectation $E_{n}$ by
$E_{n}(f)=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}f_{Q}\chi_{Q},\quad\forall f\in L_{1}(\operatorname{\mathcal{N}}),$
where $\chi_{Q}$ is the characteristic function of $Q$, and $f_{Q}=\frac{1}{|Q|}\int_{Q}f(y){\rm d}y$. Thus, $(\operatorname{\mathcal{N}}_{n})_{n\in\operatorname{\mathbb{Z}}}$ becomes a filtration with the associated conditional expectations $(E_{n})_{n\in\operatorname{\mathbb{Z}}}$. Define $f_{n}=E_{n}(f)$. Then $(f_{n})_{n\in\operatorname{\mathbb{Z}}}$ is clearly an $L_{1}$-martingale.
Set
$\operatorname{\mathcal{N}}_{c,+}=\left\{f:\operatorname{\mathbb{R}}^{d}\rightarrow\operatorname{\mathcal{M}}\cap L_{1}(\operatorname{\mathcal{M}}):f\geq 0,\ \overrightarrow{\text{supp}}\ f\text{ is compact}\right\},$
where $\overrightarrow{\text{supp}}\ f$ means the support of $f$ as an operator-valued function in $\operatorname{\mathbb{R}}^{d}$. Based on the density of $\operatorname{\mathcal{N}}_{c,+}$ in $L_{1}(\operatorname{\mathcal{N}})_{+}$, we only need to consider the functions in $\operatorname{\mathcal{N}}_{c,+}$.
The following lemma, established in [Cuculescu1971Martingales] (see also [Parcet2009Pseudo-localization]), plays an important role in the noncommutative Calderón-Zygmund decomposition.
Lemma 2.1.
$(\rm{Cuculescu})$ Let $f\in\operatorname{\mathcal{N}}_{c,+}$. For any $\lambda>0$, there exists a decreasing sequence $(q_{n}(f,\lambda))_{n\in\operatorname{\mathbb{Z}}}$ (for convenience, we write $q_{n}(f,\lambda)$ as $q_{n}$) of projections in $\operatorname{\mathcal{N}}$ satisfying
  1. (i)
    $q_{n}\in\operatorname{\mathcal{N}}_{n}$;
  2. (ii)
    $q_{n}$ commutes with $q_{n-1}f_{n}q_{n-1}$;
  3. (iii)
    $q_{n}f_{n}q_{n}\leq\lambda q_{n}$;
  4. (iv)
    The following estimate holds
    $\phi\bigg({\textbf{1}}_{\operatorname{\mathcal{N}}}-\bigwedge_{n\in\operatorname{\mathbb{Z}}}q_{n}\bigg)\leq\frac{1}{\lambda}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We see that $f_{n}$ can be bounded by any $\lambda>0$ if $n$ is small enough. Hence, for a given $\lambda>0$, there exists an integer $n_{\lambda}(f)$ such that $f_{n}\leq\lambda{\textbf{1}}_{\operatorname{\mathcal{N}}}$ when $n\leq n_{\lambda}(f)$. The decreasing sequence of projections $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ in Lemma 2.1 is in fact given by
$q_{n}=\begin{cases}{\textbf{1}}_{\operatorname{\mathcal{N}}},&\text{if}\ n\leq n_{\lambda}(f),\\ \chi_{(0,\lambda](q_{n-1}f_{n}q_{n-1})},&\text{if}\ n>n_{\lambda}(f).\end{cases}$
The following alternative representation of $q_{n}$ is more useful:
$q_{n}=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}q_{Q}\chi_{Q},$
where
$q_{Q}=\begin{cases}{\textbf{1}}_{\operatorname{\mathcal{N}}},&\text{if}\ n\leq n_{\lambda}(f),\\ \chi_{(0,\lambda](q_{\widehat{Q}}f_{Q}q_{\widehat{Q}})},&\text{if}\ n>n_{\lambda}(f).\end{cases}$
Here, $\widehat{Q}$ denotes the dyadic father of $Q$. One can easily verify that
(2.1) $q_{Q}\leq q_{\widehat{Q}},\quad q_{Q}\text{ commutes with }q_{\widehat{Q}}f_{Q}q_{\widehat{Q}},\quad q_{Q}f_{Q}q_{Q}\leq\lambda q_{Q}.$
Furthermore, let $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be a sequence of projections defined by
$p_{n}=q_{n-1}-q_{n}=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(q_{\widehat{Q}}-q_{Q})\chi_{Q}=:\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}p_{Q}\chi_{Q}.$
This gives
(2.2) $\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}={\textbf{1}}_{\operatorname{\mathcal{N}}}-q=q^{\perp}\quad\text{with}\quad q=\bigwedge_{n\in\operatorname{\mathbb{Z}}}q_{n}.$
The next result follows directly from Lemma 2.1; we omit the proofs here.
Lemma 2.2.
Let $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ and $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be the sequences of projections in Lemma 2.1. Then, the following equalities hold:
  1. (i)
    $p_{n}f_{n}q=qf_{n}p_{n}=0$, for $n\in\operatorname{\mathbb{Z}}$.
  2. (ii)
    $p_{n}f_{n\wedge k}p_{k}=p_{k}f_{n\wedge k}p_{n}=0$, for $n$, $k\in\operatorname{\mathbb{Z}}$, $n\neq k$, where $n\wedge k:=\min\left\{n,k\right\}$.
2.2. Noncommutative Calderón-Zygmund decomposition
For fixed $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$, we define the projection $\zeta_{s}$ as
(2.3) $\zeta_{s}=\bigg(\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}\bigg)^{\perp}.$
We require the following property of this projection, whose proof can be found in [Cadilhac2022Spectral].
Lemma 2.3.
For fixed $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$, we have
$\phi({\textbf{1}}_{\operatorname{\mathcal{N}}}-\zeta_{s})\leq\frac{(2s+1)^{d}}{\lambda}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Below, we construct a slightly new noncommutative Calderón-Zygmund decomposition inspired by [Cadilhac2022Spectral], and present its properties later.
Theorem 2.4.
Fix $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$. Let $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ and $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be the sequences of projections in Lemma 2.1, where $q_{n}=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}q_{Q}\chi_{Q}$, $p_{n}=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}p_{Q}\chi_{Q}$. Suppose $h\in C_{c}^{\infty}(\operatorname{\mathbb{R}}^{d})$ satisfies
$\operatorname{supp}h\subset B_{1},\quad\int_{\operatorname{\mathbb{R}}^{d}}h=1,\quad h\geq 0.$
Define $h_{r}(x)=\frac{1}{r^{d}}h(\frac{x}{r})$ and $r_{n}=2^{-n-1}\sqrt{d}$. Then, $f$ admits the decomposition:
$f=g+b=g+\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0},$
where $g$ is positive and $b$ is self-adjoint in $\operatorname{\mathcal{N}}$. More precisely, these components are defined as follows:
$g=qfq+\sum\limits_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}p_{n},$
$\widetilde{b}_{\text{d}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{d},n}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{d},n}^{Q},\quad b_{\text{d}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{d},n}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{d},n}^{Q}-\widetilde{b}_{\text{d},n}^{Q}),$
$\widetilde{b}_{\text{off}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{off},n}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{off},n}^{Q},\quad b_{\text{off}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{off},n}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q}),$
where for each $Q\in\operatorname{\mathcal{Q}}_{n}$,
$b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\quad b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},$
$\widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{r_{n}},\quad\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}.$
Proof.
[…]
$f=\sum_{n\in\widehat{\operatorname{\mathbb{Z}}}}\sum_{k\in\widehat{\operatorname{\mathbb{Z}}}}p_{n}f_{n\wedge k}p_{k}+\sum_{n\in\widehat{\operatorname{\mathbb{Z}}}}\sum_{k\in\widehat{\operatorname{\mathbb{Z}}}}p_{n}(f-f_{n\wedge k})p_{k}=:g+b.$
For the good function $g$, Lemma 2.2 implies
$g=qfq+\sum_{k\in\operatorname{\mathbb{Z}}}qf_{k}p_{k}+\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}q+\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k\in\operatorname{\mathbb{Z}}}p_{n}f_{n\wedge k}p_{k}=qfq+\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}p_{n}.$
For the bad function $b$, we consider those terms with finite indices $n$ and $k$ satisfying $n\neq k$. Since $p_{k}=q_{k-1}-q_{k}$, a direct calculation shows
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k>n}p_{n}(f-f_{n})p_{k}=\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}(f-f_{n})(q_{n}-q),$
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k<n}p_{n}(f-f_{k})p_{k}=\sum_{k\in\operatorname{\mathbb{Z}}}\sum_{n>k}p_{n}(f-f_{k})p_{k}=\sum_{k\in\operatorname{\mathbb{Z}}}(q_{k}-q)(f-f_{k})p_{k},$
which implies that
$b=\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}(f-f_{n})p_{n}+\sum_{n\in\operatorname{\mathbb{Z}}}\big(p_{n}(f-f_{n})q_{n}+q_{n}(f-f_{n})p_{n}\big)$
by using Lemma 2.2 again. Next, we insert a convolution operator to decompose $b$ further. Let $h_{r}$ and $r_{n}$ be defined as above, and let
$\left\{\begin{array}[]{ll}b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{r_{n}},\end{array}\right.\quad\text{and}\quad\left\{\begin{array}[]{ll}b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}.\end{array}\right$.
By these definitions as well as the disjointness of dyadic cubes, we can express the bad function $b$ as
$\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{d},n}^{Q}}_{\widetilde{b}_{\text{d}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{d},n}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{d},n}^{Q}-\widetilde{b}_{\text{d},n}^{Q})}_{b_{\text{d}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{d},n}^{0}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{off},n}^{Q}}_{\widetilde{b}_{\text{off}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{off},n}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q})}_{b_{\text{off}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{off},n}^{0}}.$
This yields the desired decomposition $f=g+b=g+\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0}$; the positivity of $g$ and self-adjointness of $b$ follows immediatly from their constructions. ∎
Remark 2.5.
For any $n\in\operatorname{\mathbb{Z}}$ and $Q\in\operatorname{\mathcal{Q}}_{n}$, $r_{n}$ is the radius of the circumscribed ball of the cube $Q$.
The decomposition in Theorem 2.4 admits the following properties:
Lemma 2.6.
Let $\zeta_{s}$ be the projection given as in (2.3). Then,
  1. (i)
    $\left\|g\right\|_{L_{1}(\operatorname{\mathcal{N}})}\leq\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$ and $\left\|g\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq 2^{d}\lambda$.
  2. (ii)
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}=0$ (i.e. the cancellation condition), and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$. Furthermore, $\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\leq 2\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
  3. (iii)
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{d},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
  4. (iv)
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{off},n}^{Q}=0$ (i.e. the cancellation condition), and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{off},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
  5. (v)
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{off},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq\big(\frac{2s+1}{\sqrt{d}}-2\big)r_{n}$.
Proof.
The property $(i)$ follows from Parcet’s noncommutative Calderón-Zygmund decomposition, as detailed in [Cadilhac2018Weak] or [Cadilhac2022Spectral].
Consider $(ii)$. By the definition of $b_{\text{d},n}^{Q}$ in Theorem 2.4, we have $\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}=0$ for each $Q\in\operatorname{\mathcal{Q}}_{n}$. Note that $\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n}}=1$, we get
$\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}(x){\rm d}x=\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}(y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n}}(x-y){\rm d}x\Big){\rm d}y=0,$
which establishes the cancellation condition for $\widetilde{b}_{\text{d},n}^{Q}$.
We now prove the support condition. For each $Q\in\operatorname{\mathcal{Q}}$, from the construction of $\zeta_{s}$ in (2.3), we see
(2.4) $\zeta_{s}(x)p_{Q}=p_{Q}\zeta_{s}(x)=0\quad\text{if}\quad x\in(2s+1)Q.$
Indeed, for a fixed $x\in(2s+1)Q$,
$\zeta_{s}(x)=1-\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}(x)\leq 1-p_{Q}.$
Hence,
$\zeta_{s}(x)p_{Q}=\zeta_{s}(x)(1-p_{Q})p_{Q}=0,$
and the same holds for $p_{Q}\zeta_{s}(x)$, which lead to (2.4). By $\operatorname{supp}(h_{r_{n}})\subset B_{r_{n}}$,
$\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)$ $=\zeta_{s}(x)\bigg(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{r_{n}}(y)}b_{\text{d},n}^{Q}(u)h_{r_{n}}(y-u){\rm d}u\bigg)\zeta_{s}(x)$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{r_{n}}(y)\bigcap Q}\Big(\zeta_{s}(x)p_{Q}(f-f_{n})(u)p_{Q}\zeta_{s}(x)\Big)h_{r_{n}}(y-u){\rm d}u.$
Using (2.4), it suffices to show:
(2.5) $x\in(2s+1)Q_{u,n}\quad\text{if}\quad|x-y|\leq\Big(\frac{2s+1}{\sqrt{d}}-2\Big)r_{n}.$
In fact, taking $u\in B_{r_{n}}(y)\bigcap Q$, we have
$|x-c(Q_{u,n})|\leq|x-y|+|y-u|+|u-c(Q_{u,n})|\leq\frac{2s+1}{\sqrt{d}}r_{n},$
implying (2.5).
The boundedness of $b_{\text{d},n}^{Q}$ results from the fact that $\phi(fp_{n})=\phi(f_{n}p_{n})$ and (2.2). This means,
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\leq\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\big(\phi(fp_{Q}\chi_{Q})+\phi(f_{n}p_{Q}\chi_{Q})\big)$
$=\sum_{n\in\operatorname{\mathbb{Z}}}\big(\phi(fp_{n})+\phi(f_{n}p_{n})\big)$
$\leq 2\phi(f).$
This completes the proof of $(ii)$.
The proofs of (iii)(iii)$(v)$ follow by analogous arguments and are therefore omitted. ∎
Remark 2.7.
We will fix $s=[6\sqrt{d}]$ throughout this paper. Furthermore, the projection $\zeta_{s}$ is denoted simply by $\zeta$ when $s$ is fixed.
Remark 2.8.
Recall the construction of $q_{n}$: there exists an integer $n_{\lambda}(f)$ such that $q_{n}=1_{\operatorname{\mathcal{N}}}$ when $n\leq n_{\lambda}(f)$. Without loss of generality, we may set $n_{\lambda}(f)=0$. This leads to
$\widetilde{b}_{\text{d}}=\sum_{n\geq 1}\widetilde{b}_{\text{d},n},\quad\ b_{\text{d}}^{0}=\sum_{n\geq 1}b_{\text{d},n}^{0},\quad\ \widetilde{b}_{\text{off}}=\sum_{n\geq 1}\widetilde{b}_{\text{off},n},\quad\ b_{\text{off}}^{0}=\sum_{n\geq 1}b_{\text{off},n}^{0},$
thereby simplifying the summation.
3. Two reductions
This section introduces two key lemmas (Lemma 3.1 and Lemma 3.2), which reduce the proof of the main theorem to the cases of real kernels and lacunary sequences. For their proofs, we refer the reader to [hong2023maximal].
Step 1. Let $T_{\epsilon}$ be a noncommutative Calderón-Zygmund truncated operator with a complex kernel $K$ (defined as in (1.9)). We decompose $K$ into its real and imaginary parts: $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$. The truncated operators associated with $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$ are then given by
$\operatorname{Re\,}(T_{\epsilon})f(x)=\int_{|x-y|>\epsilon}\operatorname{Re\,}(K)(x,y)f(y){\rm d}y$
and
$\operatorname{Im\,}(T_{\epsilon})f(x)=\int_{|x-y|>\epsilon}\operatorname{Im\,}(K)(x,y)f(y){\rm d}y.$
Lemma 3.1.
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.8)) whose kernel $K$ satisfies (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}$ be defined as in (1.9). Then,
  1. (i)
    Both $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$ satisfy (1.1) and (1.7).
  2. (ii)
    If $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $p_{0}\in(1,\infty)$, then, both $(\operatorname{Re\,}(T_{\epsilon}))_{\epsilon>0}$ and $(\operatorname{Im\,}(T_{\epsilon}))_{\epsilon>0}$ are of strong type $(p_{0},p_{0})$.
  3. (iii)
    If $(\operatorname{Re\,}(T_{\epsilon}))_{\epsilon>0}$ and $(\operatorname{Im\,}(T_{\epsilon}))_{\epsilon>0}$ are of weak type $(1,1)$, so is $(T_{\epsilon})_{\epsilon>0}$.
By this lemma, we may assume that the kernel $K$ is real throughout this paper.
Step 2. Let $\Phi$ be a smooth radial nonnegative function on $\operatorname{\mathbb{R}}^{d}$ such that
$\operatorname{supp}\Phi\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:1/2\leq|x|\leq 2\big\}\quad\text{and}\quad\sum_{j\in\operatorname{\mathbb{Z}}}\Phi_{j}(x)=1,\ x\in\operatorname{\mathbb{R}}^{d}\setminus\{0\},$
where $\Phi_{j}(x):=\Phi(\frac{x}{2^{j}\sqrt{d}})$. Thus, the truncated singular integrals $T_{\epsilon}f$ can be written as
$T_{\epsilon}f=\sum_{j\in\operatorname{\mathbb{Z}}}\int_{|x-y|>\epsilon}K(x,y)\Phi_{j}(x-y)f(y){\rm d}y.$
From the conditions $|x-y|>\epsilon$ and $\operatorname{supp}(\Phi_{j})\subset\{x\in\mathbb{R}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\}$, we see that for $i_{\epsilon}=\big[\log_{2}(\frac{\epsilon}{2\sqrt{d}})\big]+1$, the following equality holds (when $\epsilon>0$ is sufficiently small):
$T_{\epsilon}f(x)=\sum_{j\geq i_{\epsilon}}\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)\Phi_{j}(x-y)f(y){\rm d}y+\int_{|x-y|>\epsilon}K(x,y)\Phi_{i_{\epsilon}}(x-y)f(y){\rm d}y.$
It is convenient to introduce the following notation:
$K_{j}(x,y)=K(x,y)\Phi_{j}(x-y),$
(3.1) $S_{i}f(x)=\sum_{j\geq i}\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)f(y){\rm d}y,$
(3.2) $T_{\epsilon}^{i_{\epsilon}}f(x)=\int_{|x-y|>\epsilon}K_{i_{\epsilon}}(x,y)f(y){\rm d}y.$
Hence,
$T_{\epsilon}f(x)=S_{i_{\epsilon}}f(x)+T_{\epsilon}^{i_{\epsilon}}f(x).$
Lemma 3.2.
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.8)) whose kernel $K$ satisfies (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $T_{\epsilon}^{i_{\epsilon}}$ be defined as in (3.2). Then
  1. (i)
    $(T_{\epsilon}^{i_{\epsilon}})_{\epsilon>0}$ is of strong type $(p,p)$ for $1<p\leq\infty$.
  2. (ii)
    $(T_{\epsilon}^{i_{\epsilon}})_{\epsilon>0}$ is of weak type $(1,1)$.
The preceding two lemmas imply that Theorem 1.1 follows directly from the weak type (1,1) boundedness of $S_{i}$ given in (3.1). Thus, it remains to prove
Theorem 3.3.
Let $T$ be a noncommutative Calderón-Zygmund operator defined as in (1.8) with the real kernel $K$ satisfying (1.1) and the $L_{2}$-mean Hörmander condition (1.7). Let $S_{i}$ be defined as in (3.1). Then, the sequence of operators $(S_{i})_{i\in\operatorname{\mathbb{Z}}}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
$\left\|(S_{i}f)_{i\in\operatorname{\mathbb{Z}}}\right\|_{\Lambda_{1,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
More precisely, for any $f\in L_{1}(\operatorname{\mathcal{N}})$ and $\lambda>0$, there exists a projection $e\in\operatorname{\mathcal{N}}$ such that
$\sup_{i\in\operatorname{\mathbb{Z}}}\left\|eS_{i}fe\right\|_{L_{\infty(\operatorname{\mathcal{N}})}}\leq\lambda\quad\text{and}\quad\phi(e^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
4. The proof of Theorem 3.3
Note that every operator in a von Neumann algebra can be decomposed into a linear combination of four positive elements. By this fact and the density of $\operatorname{\mathcal{N}}_{c,+}$ in $L_{1}(\operatorname{\mathcal{N}})_{+}$, we may assume $f\in\operatorname{\mathcal{N}}_{c,+}$ without loss of generality. Now, consider a fixed $f\in\operatorname{\mathcal{N}}_{c,+}$ and $\lambda>0$. We apply Theorem 2.4 to decompose $f=g+b$. Then, by the quasi-triangle inequality, it suffices to find projections $e_{1},e_{2}\in\operatorname{\mathcal{N}}$ such that
(4.1) $\sup_{i\in\operatorname{\mathbb{Z}}}\left\|e_{1}S_{i}ge_{1}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda\quad\text{and}\quad\phi(e_{1}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
(4.2) $\sup_{i\in\operatorname{\mathbb{Z}}}\left\|e_{2}S_{i}be_{2}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda\quad\text{and}\quad\phi(e_{2}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Indeed, by setting $e=e_{1}\wedge e_{2}$, we obtain
$\left\|eS_{i}fe\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\left\|eS_{i}ge\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}+\left\|eS_{i}be\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq 2\lambda$
and
$\phi(e^{\bot})\leq\phi(e_{1}^{\bot})+\phi(e_{2}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which gives Theorem 3.3.
4.1. Estimate for the good function (4.1)
For any $i\in\operatorname{\mathbb{Z}}$, it follows from the definitions of $S_{i}$ and $T_{\epsilon}^{i_{\epsilon}}$ ((3.1) and (3.2)) that there exists $\epsilon_{i}>0$ such that
$S_{i}=S_{\epsilon_{i}}-T_{\epsilon_{i}}^{i}.$
Under the assumptions of Theorem 1.1 and Lemma 3.2 $(i)$, both $(S_{\epsilon_{i}})_{i\in\mathbb{Z}}$ and $(T_{\epsilon_{i}}^{i})_{i\in\mathbb{Z}}$ are of strong type $(p_{0},p_{0})$. Therefore, from the above equality, $(S_{i})_{i\in\mathbb{Z}}$ is also of strong type $(p_{0},p_{0})$. In addition, since $g$ is positive (by Theorem 2.4) and the kernel $K$ is real-valued (as stated in Section 3), the operator $S_{i}g$ is self-adjoint. These two facts imply that for any $i\in\operatorname{\mathbb{Z}}$, we can find a positive element $a\in\operatorname{\mathcal{N}}$ such that for any $i\in\operatorname{\mathbb{Z}}$,
$a\leq S_{i}g\leq a\quad\text{and}\quad\left\|a\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}\lesssim\left\|g\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}.$
Taking $e_{1}=\chi_{(0,\lambda]}(a)$. Then we have
$\lambda\leq-e_{1}ae_{1}\leq e_{1}S_{i}ge_{1}\leq e_{1}ae_{1}\leq\lambda.$
Finally, applying the Chebyshev inequality, the Hölder inequality, along with the boundedness of $g$ in Lemma 2.6 $(i)$, we obtain
$\phi(e_{1}^{\bot})\leq\lambda^{-p_{0}}\left\|a\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}^{p_{0}}\leq\lambda^{-p_{0}}\left\|g\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}^{p_{0}-1}\left\|g\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which yields (4.1).
4.2. Estimate for the bad function (4.2)
We now give an estimate for the bad function $b$. First, we decompose $S_{i}b$ as
$S_{i}b=\zeta^{\bot}S_{i}b\zeta^{\bot}+\zeta S_{i}b\zeta^{\bot}+\zeta^{\bot}S_{i}b\zeta+\zeta S_{i}b\zeta,$
where the projection $\zeta$ is defined as in (2.3), and we omit its subscript as stated in Remark 2.7.
For each $j\in\operatorname{\mathbb{Z}}$, define the operator $\mathcal{T}_{j}$ by
$\mathcal{T}_{j}f(x)=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)f(y){\rm d}y.$
Then, $S_{i}$ admits a representation $S_{i}=\sum\limits_{j\geq i}\mathcal{T}_{j}$. The key step is to establish the following $L_{1}$-estimate:
(4.3) $\sum_{j\in\operatorname{\mathbb{Z}}}\left\|\zeta\mathcal{T}_{j}b\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We claim that (4.3) holds, and use it to deduce (4.2). The proof of (4.3) itself will be given later.
Define the projection $\eta=\chi_{(0,\lambda]}\Big(\sum\limits_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\Big)$. It then follows that
$-\lambda\leq-\eta\sum_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\eta\leq\eta\zeta S_{i}b\zeta\eta\leq\eta\sum_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\eta\leq\lambda.$
Moreover, by the Chebyshev inequality and (4.3),
$\phi(\eta^{\perp})\leq\lambda^{-1}\sum_{j\in\operatorname{\mathbb{Z}}}\left\|\zeta\mathcal{T}_{j}b\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Now, let $e_{2}=\zeta\wedge\eta$. Combining these results with Lemma 2.3, we obtain
$\left\|e_{2}S_{i}be_{2}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\left\|\eta\zeta S_{i}b\zeta\eta\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda$
and
$\phi(e_{2})\leq\phi(\zeta^{\perp})+\phi(\eta^{\perp})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which leads to (4.2).
It remains to prove (4.3). Recall from Theorem 2.4 that the bad function $b$ has the decomposition:
$b=\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0}.$
Therefore, by the triangle inequality, it suffices to verify the following two lemmas:
Lemma 4.1.
$(i)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$; $(ii)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}b^{0}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
Lemma 4.2.
$(i)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$; $(ii)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}b_{\text{off}}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
4.2.1. The proof of Lemma 4.1
First, let us give an important lemma.
Lemma 4.3.
Fix $1\leq q<\infty$ and $j\in\operatorname{\mathbb{Z}}$. Then, for $n\geq 1$,
$\sup_{y\in\operatorname{\mathbb{R}}^{d}}\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}$ $|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x{\rm d}z$
$\lesssim 2^{-jd(q-1)}\big(H_{q}^{q}(n)+H_{q}^{q}(n+1)+2^{-qn}\big),$
where $r_{n-j}$ is defined in Theorem 2.4.
Proof.
For a fixed $z\in B_{r_{n-j}}$, we have
$\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x$ $\lesssim\int_{\operatorname{\mathbb{R}}^{d}}|K(x,y+z)-K(x,y)|^{q}|\Phi_{j}(x-y)|^{q}{\rm d}x$
$\ \ +\int_{\operatorname{\mathbb{R}}^{d}}|K(x,y+z)|^{q}|\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|^{q}{\rm d}x$
$=:A_{1}+A_{2}.$
By the mean value theorem and $z\in B_{r_{n-j}}$,
$A_{2}\lesssim\int_{E_{j,n,y,z}}|K(x,y+z)|^{q}|z|^{q}2^{-jq}{\rm d}x\lesssim 2^{-qn}\int_{E_{j,n,y,z}}|K(x,y+z)|^{q}{\rm d}x,$
where $E_{j,n,y,z}$ denotes the support of the function $x\mapsto\big(\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\big)$. We claim that there exist two constants $C_{1,d}$, $C_{2,d}$ such that
(4.4) $E_{j,n,y,z}\subseteq\big\{x\in\operatorname{\mathbb{R}}^{d}:C_{1,d}2^{j}\leq|x-(y+z)|\leq C_{2,d}2^{j}\big\}.$
To prove this, note first that $\operatorname{supp}(\Phi_{j})\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$. Hence, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, at least one of the following conditions holds:
  1. $(i)$
    $2^{j-1}\sqrt{d}\leq|x-(y+z)|\leq 2^{j+1}\sqrt{d}$,
  2. $(ii)$
    $2^{j-1}\sqrt{d}\leq|x-y|\leq 2^{j+1}\sqrt{d}$.
For case $(i)$, we have $|x-(y+z)|\approx 2^{j}$ immediately. For case $(ii)$, since $r_{n-j}\leq 2^{j-2}\sqrt{d}$ when $n\geq 1$,
$|x-(y+z)|\leq|x-y|+|z|\leq 2^{j+1}\sqrt{d}+2^{j-2}\sqrt{d},$
$|x-(y+z)|\geq|x-y|-|z|\geq 2^{j-1}\sqrt{d}-2^{j-2}\sqrt{d},$
which implyies $|x-(y+z)|\approx 2^{j}$. By the estimate for $E_{j,n,y,z}$ as above with the size condition (1.1), we obtain a bound on $A_{2}$:
$A_{2}\lesssim 2^{-jd(q-1)}\cdot 2^{-qn}.$
For $A_{1}$, the definition of $r_{n-j}$ gives
$A_{1}$ $\leq\int_{2^{j-1}\sqrt{d}\leq|x-y|\leq 2^{j+1}\sqrt{d}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x$
$=\int_{2^{n}r_{n-j}\leq|x-y|\leq 2^{n+1}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x$
$\quad+\int_{2^{n+1}r_{n-j}\leq|x-y|\leq 2^{n+2}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x.$
Finally, combining the estimates for $A_{1}$, $A_{2}$, and the $L_{2}$-mean Hörmander condition (1.7), we get
$\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{q}{\rm d}x{\rm d}z$
$\quad\leq\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\Big(\int_{2^{n}r_{n-j}\leq|x-y|\leq 2^{n+1}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x\Big){\rm d}z$
$\quad\quad+\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\Big(\int_{2^{n+1}r_{n-j}\leq|x-y|\leq 2^{n+2}r_{n-j}}|K(x,y+z)-K(x,y)|^{q}{\rm d}x\Big){\rm d}z$
$\quad\quad+2^{-jd(q-1)}\cdot 2^{-qn}$
$\quad\lesssim 2^{-jd(q-1)}\big(H_{q}^{q}(n)+H_{q}^{q}(n+1)+2^{-qn}\big),$
which completes the proof. ∎
Proof of Lemma 4.1.
$(i)$ We first observe that, for any $j\in\operatorname{\mathbb{Z}}$,
(4.5) $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta=0,\quad\text{when}\ n\leq 1.$
Indeed, for a fixed $x\in\operatorname{\mathbb{R}}^{d}$, Remark 2.7 ($s=[6\sqrt{d}]$) and Lemma 2.6 $(ii)$ yield
$\zeta(x)\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)\zeta(x)=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\zeta(x)\widetilde{b}_{\text{d},n-j}(y)\zeta(x)\Big)\chi_{\big\{|x-y|>\big(\frac{2[6\sqrt{d}]+1}{\sqrt{d}}-2\big)r_{n-j}\big\}}(y){\rm d}y.$
For $n\leq 1$, the definition of $r_{n-j}$ (see Theorem 2.4) directly implies
$2^{j+1}\sqrt{d}=2^{n+2}r_{n-j}<\Big(\frac{2[6\sqrt{d}]+1}{\sqrt{d}}-2\Big)r_{n-j}.$
Since $\operatorname{supp}(\Phi_{j})\subseteq\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$, we get
$K_{j}(x,y)=K(x,y)\Phi_{j}(x-y)=0,$
which shows that $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=0$. A similar argument works for $\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta$. Thus, we obtain (4.5)
We begin with estimating $\widetilde{b}_{\text{d}}$. By Remark 2.8 and (4.5), it suffices to establish
$\sum_{n\geq 2}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
It follows from the definition of $\widetilde{b}_{\text{d},n}$ in Theorem 2.4, the cancellation condition in Lemma 2.6 $(ii)$, and the definition of $\widetilde{b}_{\text{d},{n-j}}^{Q}$ in Theorem 2.4 that
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)$ $=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)\widetilde{b}_{\text{d},{n-j}}^{Q}(y){\rm d}y$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)b_{\text{d},n-j}^{Q}(u)h_{r_{n-j}}(y-u){\rm d}y{\rm d}u.$
By the facts that $\operatorname{supp}(h_{r_{n-j}})\subseteq B_{r_{n-j}}$, $\operatorname{supp}(b^{Q}_{\text{d},n-j})\subseteq Q$, and the Minkowski inquality, we have
$\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\frac{\left\|h\right\|_{\infty}}{(r_{n-j})^{d}}\Big(\int_{Q}\int_{B_{r_{n-j}}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y{\rm d}u\Big){\rm d}x$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}\Big(\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}(u)}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}x{\rm d}y\Big){\rm d}u.$
For a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, note that the variable $u$ in the outer integral is restricted to $Q$. Moverover, by the definition of $r_{n-j}$,
$|y-c(Q)|\leq|y-u|+|u-c(Q)|\leq r_{n-j}+r_{n-j}=r_{n-j-1}.$
This implies that whenever $u\in Q$ and $y\in B_{r_{n-j}}(u)$, we have $y\in B_{r_{n-j-1}}(c(Q))$. Therefore, let z=yc(Q)z=y-c(Q),
$\frac{1}{|B_{r_{n-j}}|}$ $\int_{B_{r_{n-j}}(u)}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}x{\rm d}y$
(4.6) $\ \lesssim\frac{1}{|B_{r_{n-j-1}}|}\int_{B_{r_{n-j-1}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}x{\rm d}z.$
By employing Lemma 4.3 with $y=c(Q)$, $q=1$, and setting $n$ to n1n-1, (4.2.1) is bounded by
$H_{1}(n-1)+H_{1}(n)+2^{-n+1}.$
Then, applying the inequality $H_{1}(n)\lesssim_{d}H_{2}(n)$, the $L_{2}$-mean Hörmander condition (1.7) and the boundedness of $b_{\text{d},n}^{Q}$ in Lemma 2.6 $(ii)$, we obtain
$\ \sum_{n\geq 2}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\ \lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\big\|b_{\text{d},n-j}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}\left(H_{1}(n-1)+H_{1}(n)+2^{-n+1}\right){\rm d}u$
$\ \lesssim\sum_{n\geq 1}\big(2H_{2}(n)+2^{-n+1}\big)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\int_{Q}\big\|b_{\text{d},m}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u$
$\ \lesssim\left\|f\right\|_{1}.$
$(ii)$ Again by Remark 2.8 and (4.5), it remains to prove:
$\sum_{n\geq 2}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$ and recall the definition of $\widetilde{b}_{\text{d},n-j}^{Q}$ in Theorem 2.4,
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}^{Q}(x)$ $=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n-j}^{Q}(y-z)h_{r_{n-j}}(z){\rm d}z{\rm d}y$
$=\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n-j}}(z)\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y+z)b_{\text{d},n-j}^{Q}(y){\rm d}y{\rm d}z.$
Moreover, by $\int_{\operatorname{\mathbb{R}}^{d}}h=1$,
$\mathcal{T}_{j}b_{\text{d},n-j}^{Q}(x)=\int_{\operatorname{\mathbb{R}}^{d}}h_{r_{n-j}}(z)\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)b_{\text{d},n-j}^{Q}(y){\rm d}y{\rm d}z.$
Noting again the supports of $h_{r_{n-j}}$, $b_{\text{d},n-j}^{Q}$ and using the Minkowski inequality, we get
$\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\ \lesssim\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{1}{|B_{r_{n-j}}|}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|\|b_{\text{d},n-j}^{Q}(y)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}z{\rm d}y\Big){\rm d}x$
$\ \lesssim\big(H_{1}(n)+H_{1}(n+1)+2^{-n}\big)\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})},$
where the last inequality follows from Lemma 4.3 with $q=1$. Finally, an application of the fact that $H_{1}(n)\lesssim_{d}H_{2}(n)$, together with the $L_{2}$-mean Hörmander condition (1.7), and the boundedness of $b_{\mathrm{d},n}^{Q}$ in Lemma 2.6 $(ii)$, yield
$\sum_{n\geq 2}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\leq\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\left(H_{1}(n)+H_{1}(n+1)+2^{-n}\right)$
$\quad\lesssim\sum_{n\geq 2}\left(2H_{2}(n)+2^{-n}\right)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\|b_{\text{d},m}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which completes the argument. ∎
4.2.2. The proof of Lemma 4.2
We first give a useful lemma.
Lemma 4.4.
For a fixed $j\in\operatorname{\mathbb{Z}}$, the following estimates hold for any $Q\in\operatorname{\mathcal{Q}}_{n-j}$ when $n\geq 2$:
  1. (i)
    $\frac{1}{|B_{r_{n-j-1}}|}\int_{\operatorname{\mathbb{R}}^{d}}$ $\Big\|\int_{Q}\Big(\int_{B_{r_{n-j-1}}(c(Q))}\left|K_{j}(x,y)-K_{j}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
    $\quad\quad\quad\quad\quad\quad\quad\lesssim\big((H_{2}^{2}(n-1)+H_{2}^{2}(n)+2^{-2(n-1)})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\big)^{\frac{1}{2}};$
  2. (ii)
    $\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}$ $\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
    $\quad\quad\quad\quad\quad\quad\lesssim\left((H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}},$
where $r_{n-j}$ is defined in Theorem 2.4.
Proof.
We first prove $(ii)$. For $n\geq 2$ and a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, using the same argument as in Lemma 4.3, we can find constants $D_{1,d}$ and $D_{2,d}$ such that the function
$x\mapsto\int_{Q}\Big({\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z}\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y$
is supported on the set
$E_{j}^{Q}:=\big\{x\in\operatorname{\mathbb{R}}^{d}:D_{1,d}2^{j}\leq\text{dist}(x,Q)\leq D_{2,d}2^{j}\big\},$
where $\text{dist}(x,Q):=\inf\left\{|x-y|:y\in Q\right\}$. Then, by the Hölder inequality and the Cauchy-Schwarz inequality, the left-hand side of the inequality $(ii)$ is bounded by $B_{1}\cdot B_{2}$, where $B_{1}$ is defined as
$\frac{1}{|B_{r_{n-j}}|}\bigg(\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x\bigg)^{\frac{1}{2}}$
and $B_{2}$ is defined as
$\Big(\int_{\operatorname{\mathbb{R}}^{d}}\big\|p_{Q}\chi_{E_{j}^{Q}}(x)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x\Big)^{\frac{1}{2}}.$
For $B_{1}$, applying the Cauchy-Schwarz inequality to the inner integral for $z$,
$(B_{1})^{2}$ $\leq\int_{Q}\Big(\frac{1}{|B_{r_{n-j}}|}\int_{B_{r_{n-j}}}\int_{\operatorname{\mathbb{R}}^{d}}|K_{j}(x,y+z)-K_{j}(x,y)|^{2}{\rm d}x{\rm d}z\Big)\tau\big(p_{Q}f(y)p_{Q}\big){\rm d}y$
$\lesssim 2^{-jd}\big(H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n}\big)\phi(fp_{Q}\chi_{Q}),$
where the last inequality follows from Lemma 4.3 with $q=2$. For $B_{2}$, the construction of $E_{j}^{Q}$ implies that
$(B_{2})^{2}\leq 2^{jd}\tau(p_{Q}).$
Combining the estimates for $B_{1}$ and $B_{2}$, we conclude that
$\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}(\operatorname{\mathcal{M}})}}^{\frac{1}{2}}{\rm d}x$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\lesssim\left((H_{2}^{2}(n)+H_{2}^{2}(n+1)+2^{-2n})\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}},$
which completes the proof of $(ii)$.
For $(i)$, we note that there also exist two constants $D_{1,d}^{\prime}$, $D_{2,d}^{\prime}$ such that the support of the function
$x\mapsto\int_{Q}\Big(\int_{B_{r_{n-j-1}(c(Q))}}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u$
is contained in the set
$\big\{x\in\operatorname{\mathbb{R}}^{d}:D_{1,d}^{\prime}2^{j}\leq|x-c(Q)|\leq D_{2,d}^{\prime}2^{j}\big\}.$
The remaining argument parallels that for $(ii)$ and is therefore omitted. ∎
Proof of Lemma 4.2.
$(i)$ As in the proof for the diagonal part $b_{\text{d}}$, we state the following equality, whose proof is similar to that of (4.5): for any $j\in\operatorname{\mathbb{Z}}$,
(4.7) $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta=\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta=0,\quad\text{when}\ n\leq 1.$
Combining this with Remark 2.8, it suffices to prove
$\sum_{n\geq 2}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We start to estimate the left-hand side. The definition of $\widetilde{b}_{\text{off},n-j}$ in Theorem 2.4 and the cancellation condition in Lemma 2.6 $(iv)$ give:
$\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\quad\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x.$
Next, recall the definition of $\widetilde{b}_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{r_{n}}$, where $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Accordingly, we decompose $\sum\limits_{n\geq 2}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ into four terms:
$F_{i}:=\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q,i}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, with
$\widetilde{b}_{\text{off},n-j}^{Q,1}:=p_{Q}\big((f\chi_{Q})*h_{r_{n-j}}\big)q_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,2}:=p_{Q}\big((f_{Q}\chi_{Q})*h_{r_{n-j}}\big)q_{Q},$
$\widetilde{b}_{\text{off},n-j}^{Q,3}:=q_{Q}\big((f\chi_{Q})*h_{r_{n-j}}\big)p_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,4}:=q_{Q}\big((f_{Q}\chi_{Q})*h_{r_{n-j}}\big)p_{Q}.$
We provide a detailed proof only for $F_{1}$; analogous arguments apply to $F_{2}$$F_{4}$ under the fact that $\phi(f_{n}p_{n})=\phi(fp_{n})$.
To estimate $F_{1}$. We first observe that by the supports of $h_{r_{n-j}}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm in $F_{1}$ equals
$\Big\|\int_{Q}\Big(\int_{|u-y|\leq r_{n-j}}(K_{j}(x,y)-K_{j}(x,c(Q)))h_{r_{n-j}}(y-u){\rm d}y\Big)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
By using Lemma A.3 (in Appendix A), the above estimate is bounded by a product
$\Big\|\Big(\int_{Q}\Big(\int_{B_{r_{n-j}}(u)}\left|K_{i}(x,y)-K_{i}(x,c(Q))\right||h_{r_{n-j}}(y-u)|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big)^{{}^{\frac{1}{2}}}\Big\|_{L_{1}(\operatorname{\mathcal{M}})}$
$\cdot\Big\|\Big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}.$
For the second term, an application of (2.1) shows that
$\Big\|\Big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\lesssim(|Q|\lambda)^{\frac{1}{2}}.$
Hence,
$F_{1}\lesssim$ $\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}(|Q|\lambda)^{\frac{1}{2}}$
(4.8) $\frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}(u)}\left|K_{i}(x,y)-K_{i}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x.$
The same argument as in (4.2.1) allows us to majorize (4.8) by
$\frac{1}{|B_{r_{n-j-1}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j-1}}(c(Q))}\left|K_{j}(x,y)-K_{j}(x,c(Q))\right|{\rm d}y\Big)^{2}p_{Q}f(u)p_{Q}{\rm d}u\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x.$
By Lemma 4.4 $(i)$ and the $L_{2}$-mean Hörmander condition (1.7), we obtain
$F_{1}$ $\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\left(\big(H_{2}^{2}(n-1)+H_{2}^{2}(n)+2^{-2(n-1)}\big)\tau(p_{Q})\phi(fp_{Q}\chi_{Q})\right)^{\frac{1}{2}}(|Q|\lambda)^{\frac{1}{2}}$
$\lesssim\sum_{n\geq 2}\left(2H_{2}(n-1)+2^{-n+1}\right)\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}.$
Finally, applying the Cauchy-Schwarz inequality twice, we estimate $F_{1}$ as follows:
$F_{1}\lesssim\Big(\sum_{m=1}^{\infty}\lambda\phi(p_{m})\Big)^{\frac{1}{2}}\Big(\sum_{m=1}^{\infty}\phi(fp_{m})\Big)^{\frac{1}{2}}=\big(\lambda\phi(1-q)\big)^{\frac{1}{2}}\big(\phi(f(1-q))\big)^{\frac{1}{2}}$ $\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
$(ii)$ We now consider $b_{\text{off}}^{0}$. Combining (4.7) with Remark 2.8, it suffices to prove
$\sum_{n\geq 2}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
First, a direct calculation gives
$\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)b_{\text{off},n-j}^{0}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x.$
As in the proof of Lemma 4.1 $(ii)$, we bound the noncommutative $L_{1}$-norm on the right-hand side by
(4.9) $\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\Big\|\int_{Q}\int_{B_{r_{n-j}}}\big(K_{j}(x,y)-K_{j}(x,y+z)\big)h_{r_{n-j}}(z)b^{Q}_{\text{off},n-j}(y){\rm d}z{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
Now, recall the definition of $b_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Substituting this into (4.9), we therefore decompose $\sum\limits_{n\geq 2}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ into four terms:
$G_{i}:=$ $\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}$
$\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{B_{r_{n-j}}}\int_{Q}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)b_{\text{off},n-j}^{Q,i}(y,z){\rm d}y{\rm d}z\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, with
$b_{\text{off},n-j}^{Q,1}(y,z)=p_{Q}(f\chi_{Q})(y)h_{r_{n-j}}(z)q_{Q},\quad b_{\text{off},n-j}^{Q,2}(y,z)=p_{Q}(f_{Q}\chi_{Q})(y)h_{r_{n-j}}(z)q_{Q},$
$b_{\text{off},n-j}^{Q,3}(y,z)=q_{Q}(f\chi_{Q})(y)h_{r_{n-j}}(z)p_{Q},\quad b_{\text{off},n-j}^{Q,4}(y,z)=q_{Q}(f_{Q}\chi_{Q})(y)h_{r_{n-j}}(z)p_{Q}.$
Again, in the following we only estimate $G_{1}$ in detail; similar arguments apply to $G_{2}$$G_{4}$ by using $\phi(f_{n}p_{n})=\phi(fp_{n})$.
Consider $G_{1}$. It follows from Lemma A.3 (in Appendix A), (2.1), Lemma 4.4 $(ii)$, and the $L_{2}$-mean Hörmander condition (1.7) that
$G_{1}$ $\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}(|Q|\lambda)^{\frac{1}{2}}$
$\ \ \ \ \frac{1}{|B_{r_{n-j}}|}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\Big(\int_{B_{r_{n-j}}}\left|K_{i}(x,y+z)-K_{i}(x,y)\right|{\rm d}z\Big)^{2}p_{Q}f(y)p_{Q}{\rm d}y\Big\|_{L_{\frac{1}{2}}(\operatorname{\mathcal{M}})}^{\frac{1}{2}}{\rm d}x$
$\lesssim\sum_{n\geq 2}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\big(2H_{2}(n)+2^{-n}\big)\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big(\tau(p_{Q})\phi(fp_{Q}\chi_{Q})|Q|\lambda\big)^{\frac{1}{2}}.$
Then, the remaining part of the proof follows the same method as used for $F_{1}$. This completes the estimate for $G_{1}$ and the proof of Lemma 4.2. ∎
Appendix A Noncommutative $L_{p}$ spaces
This part consists of three subsections: noncommutative $L_{p}$ spaces, noncommutative maximal norms, and noncommutative square functions. All these content are standard in noncommutative harmonic analysis, we refer to see [Cuculescu1971Martingales, Junge2002Doob, Parcet2009Pseudo-localization, Pisier1998noncommutative, PisierQuanhua2003] for more details.
A.1. Noncommutative $L_{p}$ spaces
[…]
Let $0<p<\infty$. For any $x\in\operatorname{\mathcal{S}}$, the operator $|x|^{p}\in\operatorname{\mathcal{S}}_{+}$ (meaning $\tau(|x|^{p})<\infty$) , where $|x|=(x^{*}x)^{\frac{1}{2}}$ is the modulus of $x$. We define
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}})}=\big(\tau(|x|^{p})\big)^{\frac{1}{p}},\quad x\in\operatorname{\mathcal{S}}.$
The noncommutative $L_{p}$ space associated with $(\operatorname{\mathcal{M}},\tau)$, denoted by $L_{p}(\operatorname{\mathcal{M}},\tau)$ or simply $L_{p}(\operatorname{\mathcal{M}})$, is defined as the completion of the space $(\operatorname{\mathcal{S}},\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}})})$. For convenience, we set $L_{\infty}(\operatorname{\mathcal{M}})=\operatorname{\mathcal{M}}$ equipped with the operator norm $\left\|\cdot\right\|_{\operatorname{\mathcal{M}}}$. We refer the reader to [PisierQuanhua2003] for more information regarding noncommutative $L_{p}$ spaces.
A.2. Noncommutative maximal norms
The generalization of the maximal function to noncommutative spaces faces a challenge: operators in von Neumann algebras are not directly comparable. Remarkably, Pisier [Pisier1998noncommutative] and Junge [Junge2002Doob] overcame this difficulty by directly defining the noncommutative maximal norm.
Definition A.1.
For $1\leq p\leq\infty$, we define $L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ as the space of all sequences $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}$ in $L_{p}(\operatorname{\mathcal{M}})$ which admit a factorization of the following form: there exist $a$, $b\in L_{2p}(\operatorname{\mathcal{M}})$ and a bounded sequence $y=(y_{n})_{n\in\operatorname{\mathbb{Z}}}$ in $L_{\infty}(\operatorname{\mathcal{M}})$ such that
$x_{n}=ay_{n}b,\quad n\in\operatorname{\mathbb{Z}}.$
The norm of $x\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ is defined as
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf_{x_{n}=ay_{n}b}\big\{\left\|a\right\|_{L_{2p}(\operatorname{\mathcal{M}})}\sup_{n\in\operatorname{\mathbb{Z}}}\left\|y_{n}\right\|_{L_{\infty}(\operatorname{\mathcal{M}})}\left\|b\right\|_{L_{2p}(\operatorname{\mathcal{M}})}\big\},$
where the infimum is taken over all factorizations of $x$ as above.
One can check that $L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ is a Banach space equipped with the noncommutative maximal norm $\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}$. A positive sequence $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ if and only if there exists a positive element $a\in L_{p}(\operatorname{\mathcal{M}})$ such that
$0<x_{n}\leq a,\quad\forall n\in\operatorname{\mathbb{Z}}.$
Moreover, its norm is given by
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf\big\{\left\|a\right\|_{L_{p}(\operatorname{\mathcal{M}})}:0<x_{n}\leq a,\ \forall n\in\operatorname{\mathbb{Z}}\big\}.$
Similarly, when $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}$ is a sequence of self-adjoint operators in $L_{p}(\operatorname{\mathcal{M}})$, $x\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ if and only if there exists a positive element $a\in L_{p}(\operatorname{\mathcal{M}})$ such that
$-a\leq x_{n}\leq a,\quad\forall n\in\operatorname{\mathbb{Z}}.$
In this case,
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf\big\{\left\|a\right\|_{L_{p}(\operatorname{\mathcal{M}})}:-a\leq x_{n}\leq a,\ a>0,\ \forall n\in\operatorname{\mathbb{Z}}\big\}.$
The concept of the weak maximal norm is also essential in this paper.
Definition A.2.
For $1\leq p<\infty$ and $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}\subset L_{p}(\operatorname{\mathcal{M}})$, we define
$\left\|x\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\lambda\inf_{e\in\operatorname{\mathcal{P}}(\operatorname{\mathcal{M}})}\big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:\left\|ex_{n}e\right\|_{L_{\infty}(\operatorname{\mathcal{M}})}\leq\lambda,\ \forall n\in\operatorname{\mathbb{Z}}\big\},$
[…]
The space $\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})$ is a quasi-Banach space. For a positive sequence $x=(x_{n})\subset L_{p}(\operatorname{\mathcal{M}})$, the weak norm also admits the equivalent expression:
$\left\|x\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\lambda\inf_{e\in\operatorname{\mathcal{P}}(\operatorname{\mathcal{M}})}\Big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:0\leq ex_{n}e\leq\lambda,\ \forall n\in\operatorname{\mathbb{Z}}\Big\}.$
A.3. Noncommutative square functions
Consider a measure space $(\Omega,\mu)$ and an operator-valued function $f$. Define
$\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}=\Big\|\Big(\int_{\Omega}f^{*}(t)f(t){\rm d}\mu(t)\Big)^{\frac{1}{2}}\Big\|_{L_{p}(\operatorname{\mathcal{M}})}.$
Then, the column space $L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))$ is defined as the space of all such functions $f$ for which $\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}<\infty$. The column space is the foundation for the following Hölder inequality, which will be useful later. We omit the symbol $\mu$ when no ambiguity arises.
Lemma A.3.
Let $0<p,q,r\leq\infty$ be such that $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$. Then for any $f\in L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$ and $g\in L_{q}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$,
$\Big\|\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega\Big\|_{L_{r}(\operatorname{\mathcal{M}})}\leq\Big\|\Big(\int_{\Sigma}|f(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}\Big\|_{L_{p}(\operatorname{\mathcal{M}})}\Big\|\Big(\int_{\Sigma}|g(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}\Big\|_{L_{q}({\mathcal{M}})}.$
We refer the reader to [PisierQuanhua2003] for more information about Hilbert valued operator spaces.