### v2

# Noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators with a class of $L_{1}$-integral conditions

Abstract.
We construct a slightly new noncommutative Calderón-Zygmund decomposition by further splitting the bad function. Using this tool, we prove the weak type (1,1) boundedness of noncommutative Calderón-Zygmund operators under a class of $L_{1}$-integral conditions, which are close to $L_{1}$-Dini conditions.
Key words and phrases:
Noncommutative Calderón-Zygmund decomposition, Noncommutative Calderón-Zygmund operators
This work is supported by National Natural Science Foundation of China (No. 12322107, No. 12271124, and No. W2441002) and Heilongjiang Provincial Natural Science Foundation of China (No. YQ2022A005).
1. Introduction
Calderón-Zygmund operators, a well-known class of singular integral operators, were introduced by Calderón and Zygmund [3] in their seminal 1952 work. The kernel of such an operator satisfies specific conditions, including the size and regularity conditions, the latter of which was refined by Hörmander [8] in 1960.
The standard nonconvolution type Calderón-Zygmund operator is defined as
$Tf(x)=\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)f(y){\rm d}y,\quad x\notin\operatorname{supp}f,$
where $K:\operatorname{\mathbb{R}}^{d}\times\operatorname{\mathbb{R}}^{d}\setminus\left\{(x,x):x\in\operatorname{\mathbb{R}}^{d}\right\}\rightarrow\operatorname{\mathbb{C}}$ is a standard nonconvolution type Calderón-Zygmund kernel, which is a binary function satisfying
  • Size condition: there exists $C_{1}>0$ such that
    (1.1) $|K(x,y)|\leq\frac{C_{1}}{|x-y|^{d}};$
  • Lipschitz regularity condition: there exists $\gamma\in[0,1]$ and $C_{2}>0$ such that
    (1.2) $|K(x,y)-K(x,z)|+|K(y,x)-K(z,x)|\leq\frac{C_{2}|y-z|^{\gamma}}{|x-y|^{d+\gamma}},\quad\text{if}\ |x-y|\geq 2|y-z|,$
where $x$, $y$, $z\in\operatorname{\mathbb{R}}^{d}$. The Lipschitz regularity condition (1.2) can be relaxed to certain variants of smoothness conditions (see [6, 20, 21]). In the following, all Calderón-Zygmund kernels under consideration are of nonconvolution type.
Inspired by operator algebras, harmonic analysis, noncommutative geometry, and quantum probability (see [1, 4, 10, 11, 13, 14, 15, 16, 17]), noncommutative harmonic analysis has become an exciting field in recent years. However, generalizing the Calderón-Zygmund decomposition, a key technique in harmonic analysis first introduced in [3], to the noncommutative setting is challenging.
In 2009, Parcet [17] rigorously constructed a kind of noncommutative Calderón-Zygmund decomposition by using Cuculescu’s maximal weak type $(1,1)$ estimates for noncommutative martingales [5]. With this tool, Parcet established the weak type $(1,1)$ estimates for operator-valued Calderón-Zygmund singular integrals with the kernel satisfying the Lipschitz regularity condition (1.2) (Cadilhac later gave a shorter proof in [1]). Notably, the Hörmander condition
(1.3) $\sup_{y\in\operatorname{\mathbb{R}}^{d},|v|>0}\int_{|x-y|\geq 2|v|}|K(x,y+v)-K(x,y)|{\rm d}x<\infty,$
which is weaker than (1.2), already guarantees the weak type $(1,1)$ boundedness in classical Calderón-Zygmund theory. It is therefore natural to ask whether the same still holds in the noncommutative setting? This problem remains open.
A step forward was made in 2022 by Cadilhac et al. [2]. They developed a more effective noncommutative Calderón-Zygmund decomposition without the off-diagonal term of the good function, which allowed them to prove the noncommutative weak type $(1,1)$ estimates for Calderón-Zygmund operators if the kernel $K$ satisfies
(1.4) $\sum_{m\geq 1}\sup_{{\begin{subarray}{c}y\in Q\\ Q\ \text{dyadic}\end{subarray}}}\Big(2^{md}\ell(Q)^{d}\int_{2^{m}\ell(Q)\leq|x-c(Q)|\leq 2^{m+1}\ell(Q)}|K(x,y)-K(x,c(Q))|^{2}{\rm d}x\Big)^{\frac{1}{2}}<\infty,$
where $\ell(Q)$ and $c(Q)$ stand for the length and the center of the dyadic cube $Q$, respectively.
More generally, for $1\leq q\leq\infty$, we say that a kernel $K$ satisfies the $L_{q}$-Dini condition if
(1.5) $\sum_{m\geq 1}\delta_{q}(m)<\infty,$
where
$\delta_{q}(m):=\sup_{{\begin{subarray}{c}y\in\operatorname{\mathbb{R}}^{d},\ R>0\\ |v|<R\end{subarray}}}$ $\Big((2^{m}R)^{d(q-1)}\int_{2^{m}R\leq|x-y|\leq 2^{m+1}R}|K(x,y+v)-K(x,y)|^{q}{\rm d}x\Big)^{\frac{1}{q}},$
for $m\geq 1$. The fact that $\delta_{1}(m)\lesssim_{d}\delta_{2}(m)$ shows that the $L_{2}$-Dini condition ((1.5) with $q=2$) is stronger than the $L_{1}$-Dini condition ((1.5) with $q=1$), but weaker than the Lipschitz regularity condition (1.2). Note that the $L_{1}$-Dini condition ((1.5) with $q=1$) corresponds to the Hörmander condition (1.3).
Using the noncommutative Calderón-Zygmund decomposition in [2], Hong, the first author, and Xu [7] established the weak type (1,1) boundedness for the noncommutative maximal truncated Calderón-Zygmund operators under the $L_{2}$-Dini condition ((1.5) with $q=2$). Further results on weak (1,1) boundedness theory for the noncommutative Calderón-Zygmund operators can be found in [2, 12, 14].
Since it remains open whether the $L_{1}$-Dini condition ((1.5) with $q=1$) guarantees the weak type $(1,1)$ boundedness, a natural intermediate problem is whether we can lower the exponent from $q=2$ to some $q$ satisfying $1\leq q<2$? This also appears to be quite challenging. In this paper, we focus on a class of $L_{1}$-integral conditions, which is close to the $L_{1}$-Dini condition ((1.5) with $q=1$). For a dyadic cube $Q$, we define the kernel average by
$\mathcal{K}_{Q}(x,y):=\frac{1}{|Q|}\int_{|v|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+v)-K(x,y)|{\rm d}v,$
and set
$\mathcal{H}(m):=\sup_{Q\ \text{dyadic}}\int_{2^{m}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{m+1}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}_{Q}(x,y){\rm d}x.$
We then say that the kernel satisfies the $L_{1}$-integral condition if
(1.6) $\sum_{m\geq 1}\mathcal{H}(m)<\infty.$
Compared with the $L_{1}$-Dini condition ((1.5) with $q=1$), the $L_{1}$-integral condition (1.6) replaces the supremum over $v$ by an average over the ball $\{v:|v|\leq\frac{\sqrt{d}}{2}\ell(Q)\}$, while the supremum over $y$ is taken inside the integral of $x$. The purpose of this paper is to show that this $L_{1}$-integral condition (1.6) is already sufficient for the weak type (1,1) boundedness of the noncommutative Calderón-Zygmund operators.
[…]
(1.7) $Tf(x)=\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)f(y){\rm d}y,\quad x\notin\overrightarrow{\operatorname{supp}}\ f,$
where $f\in L_{1}(\operatorname{\mathcal{N}})\bigcap L_{\infty}(\operatorname{\mathcal{N}})$ is compactly supported and measurable. Here, $\overrightarrow{\text{supp}}\ f$ denotes the support of $f$ as an operator-valued function in $\operatorname{\mathbb{R}}^{d}$ rather than the support projection as an element in a von Neumann algebra. For any $\epsilon>0$, the associated truncated singular integral $T_{\epsilon}f$ is defined as
(1.8) $T_{\epsilon}f(x)=\int_{|x-y|>\epsilon}K(x,y)f(y){\rm d}y.$
The main result is stated as follows.
Theorem 1.1.
Suppose $T$ is a Calderón-Zygmund operator defined as in (1.7) associated with a kernel satisfying (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}$ be defined as in (1.8). Suppose $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $1<p_{0}<\infty$, that is, for any $f\in L_{p_{0}}(\operatorname{\mathcal{N}})$,
$\left\|(T_{\epsilon}f)_{\epsilon>0}\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}.$
Then, $(T_{\epsilon})_{\epsilon>0}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
$\left\|(T_{\epsilon}f)_{\epsilon>0}\right\|_{\Lambda_{1,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We refer to see Appendix A for the definitions of the noncommutative maximal norm $\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{N}};\ell_{\infty})}$ and the noncommutative weak maximal norm $\left\|\cdot\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}$.
Our main strategy in this paper is to develop a refined noncommutative Calderón-Zygmund decomposition, along with related techniques, adapted to kernels satisfying the $L_{1}$-integral condition (1.6). Existing methods are inadequate for this purpose. Parcet’s decomposition [17] requires the stronger Lipschitz regularity condition (1.2) and may not be directly applied to the noncommutative maximal operators. Meanwhile, the decomposition by Cadilhac et al. [2] can only deal with the $L_{2}$-Dini condition ((1.5) with $q=2$) and therefore does not apply to our setting. To address this gap, we introduce a slightly new noncommutative Calderón-Zygmund decomposition achieved by further splitting of the bad function into a convolution part and a remainder part. In addition, we avoid using the Hölder inequality for the noncommutative square functions but instead employ more explicit factorizations for such terms.
This paper is organized as follows. We begin by presenting our key tool: the refined noncommutative Calderón-Zygmund decomposition (Theorem 2.4) in Section 2. Next, Section 3 starts the proof of the main result (Theorem 1.1), reducing it to proving the weak type (1,1) estimates for the lacunary sequences with real kernels (Theorem 3.3). With this reduction established, Section 4 is devoted to proving Theorem 3.3. Finally, some related background concepts are included in Appendix A.
Now, we conclude this section by listing the notation required for our later analysis.
Notation
  • $\operatorname{\mathcal{Q}}_{n}$: the set of all dyadic cubes of side length $2^{-n}$, $n\in\operatorname{\mathbb{Z}}$.
  • $Q_{x,n}$: the unique cube in $\operatorname{\mathcal{Q}}_{n}$ containing $x\in\operatorname{\mathbb{R}}^{d}$.
  • $\operatorname{\mathcal{Q}}$: the set of all standard dyadic cubes in $\operatorname{\mathbb{R}}^{d}$, that is, $\operatorname{\mathcal{Q}}=\bigcup\limits_{n}\operatorname{\mathcal{Q}}_{n}$.
  • $c(Q)$: the center of the cube $Q$.
  • lQlQ: the concentric cube sharing the center of $Q$ such that its length is $l$ times the length of $Q$.
  • $B_{r}(x)$: the ball with $x\in\operatorname{\mathbb{R}}^{d}$ as its center and $r>0$ as its radius, and the center $x$ will be omitted when it is the origin.
  • $|E|$: the volume of set $E\subset\operatorname{\mathbb{R}}^{d}$.
  • $[k]$: the integer part of $k$.
2. Noncommutative Calderón-Zygmund decomposition
The Calderón-Zygmund decomposition is a fundamental tool for studying the boundedness of singular integral operators. When transferring to the von Neumann algebraic setting, its noncommutative counterpart is primarily based on Cuculescu’s martingale projections [5]. In this section, we first review Cuculescu’s construction, and then use it to develop our refined version of the noncommutative Calderón-Zygmund decomposition. Background on noncommutative $L_{p}$ spaces and maximal norms can be found in Appendix A.
2.1. Cuculescu’s construction
Let $\operatorname{\mathcal{N}}=L_{\infty}(\operatorname{\mathbb{R}}^{d})\bar{\otimes}\operatorname{\mathcal{M}}$ be the von Neumann algebra tensor product, equipped with the tensor trace $\phi=\int_{\operatorname{\mathbb{R}}^{d}}\otimes\tau$. For $n\in\mathbb{Z}$, let $\sigma_{n}$ be the $\sigma$-algebra generated by $\operatorname{\mathcal{Q}}_{n}$, and let $\operatorname{\mathcal{N}}_{n}=L_{\infty}(\operatorname{\mathbb{R}}^{d},\sigma_{n},dx)\bar{\otimes}\operatorname{\mathcal{M}}$, which is a von Neumann subalgebra of $\operatorname{\mathcal{N}}$. Then, $(\operatorname{\mathcal{N}}_{n})_{n\in\mathbb{Z}}$ forms a sequence of increasing von Neumann subalgebras of $\operatorname{\mathcal{N}}$ whose union is weak-* dense in $\operatorname{\mathcal{N}}$.
Let $E_{n}$ be a conditional expectation from $\operatorname{\mathcal{N}}$ onto $\operatorname{\mathcal{N}}_{n}$. A sequence $(a_{n})_{n\in\operatorname{\mathbb{Z}}}\subset\operatorname{\mathcal{N}}$ is said to be a martingale if $E_{n}(a_{n+1})=a_{n}$. For $1\leq q\leq\infty$, if in addition $(a_{n})_{n\in\operatorname{\mathbb{Z}}}\subset L_{q}(\operatorname{\mathcal{N}})$, then $(a_{n})_{n\in\operatorname{\mathbb{Z}}}$ is called an $L_{q}$-martingale. In this paper, we define the conditional expectation $E_{n}$ by
$E_{n}(f)=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}f_{Q}\chi_{Q},\quad\forall f\in L_{1}(\operatorname{\mathcal{N}}),$
where $\chi_{Q}$ is the characteristic function of $Q$, and $f_{Q}=\frac{1}{|Q|}\int_{Q}f(y){\rm d}y$. Thus, $(\operatorname{\mathcal{N}}_{n})_{n\in\operatorname{\mathbb{Z}}}$ becomes a filtration with the associated conditional expectations $(E_{n})_{n\in\operatorname{\mathbb{Z}}}$. Define $f_{n}=E_{n}(f)$. Then $(f_{n})_{n\in\operatorname{\mathbb{Z}}}$ is clearly an $L_{1}$-martingale.
Set
$\operatorname{\mathcal{N}}_{c,+}=\left\{f:\operatorname{\mathbb{R}}^{d}\rightarrow\operatorname{\mathcal{M}}\cap L_{1}(\operatorname{\mathcal{M}}):f\geq 0,\ \overrightarrow{\text{supp}}\ f\text{ is compact}\right\},$
where $\overrightarrow{\text{supp}}\ f$ means the support of $f$ as an operator-valued function in $\operatorname{\mathbb{R}}^{d}$. Based on the density of $\operatorname{\mathcal{N}}_{c,+}$ in $L_{1}(\operatorname{\mathcal{N}})_{+}$, we only need to consider the functions in $\operatorname{\mathcal{N}}_{c,+}$.
The following lemma, established in [5] (see also [17]), plays an important role in the noncommutative Calderón-Zygmund decomposition.
Lemma 2.1.
$(\rm{Cuculescu})$ Let $f\in\operatorname{\mathcal{N}}_{c,+}$. For any $\lambda>0$, there exists a decreasing sequence $(q_{n}(f,\lambda))_{n\in\operatorname{\mathbb{Z}}}$ (for convenience, we write $q_{n}(f,\lambda)$ as $q_{n}$) of projections in $\operatorname{\mathcal{N}}$ satisfying
  1. (i)
    $q_{n}\in\operatorname{\mathcal{N}}_{n}$;
  2. (ii)
    $q_{n}$ commutes with $q_{n-1}f_{n}q_{n-1}$;
  3. (iii)
    $q_{n}f_{n}q_{n}\leq\lambda q_{n}$;
  4. (iv)
    The following estimate holds
    $\phi\bigg({\textbf{1}}_{\operatorname{\mathcal{N}}}-\bigwedge_{n\in\operatorname{\mathbb{Z}}}q_{n}\bigg)\leq\frac{1}{\lambda}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We see that $f_{n}$ can be bounded by any $\lambda>0$ if $n$ is small enough. Hence, for a given $\lambda>0$, there exists an integer $n_{\lambda}(f)$ such that $f_{n}\leq\lambda{\textbf{1}}_{\operatorname{\mathcal{N}}}$ when $n\leq n_{\lambda}(f)$. The decreasing sequence of projections $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ in Lemma 2.1 is in fact given by
$q_{n}=\begin{cases}{\textbf{1}}_{\operatorname{\mathcal{N}}},&\text{if}\ n\leq n_{\lambda}(f),\\ \chi_{(0,\lambda](q_{n-1}f_{n}q_{n-1})},&\text{if}\ n>n_{\lambda}(f).\end{cases}$
The following alternative representation of $q_{n}$ is more useful:
$q_{n}=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}q_{Q}\chi_{Q},$
where
$q_{Q}=\begin{cases}{\textbf{1}}_{\operatorname{\mathcal{N}}},&\text{if}\ n\leq n_{\lambda}(f),\\ \chi_{(0,\lambda](q_{\widehat{Q}}f_{Q}q_{\widehat{Q}})},&\text{if}\ n>n_{\lambda}(f).\end{cases}$
Here, $\widehat{Q}$ denotes the dyadic father of $Q$. One can easily verify that
(2.1) $q_{Q}\leq q_{\widehat{Q}},\quad q_{Q}\text{ commutes with }q_{\widehat{Q}}f_{Q}q_{\widehat{Q}},\quad q_{Q}f_{Q}q_{Q}\leq\lambda q_{Q}.$
Furthermore, let $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be a sequence of projections defined by
$p_{n}=q_{n-1}-q_{n}=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(q_{\widehat{Q}}-q_{Q})\chi_{Q}=:\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}p_{Q}\chi_{Q}.$
This gives
(2.2) $\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}={\textbf{1}}_{\operatorname{\mathcal{N}}}-q=q^{\perp}\quad\text{with}\quad q=\bigwedge_{n\in\operatorname{\mathbb{Z}}}q_{n}.$
The next result follows directly from Lemma 2.1; we omit the proofs here.
Lemma 2.2.
Let $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ and $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be the sequences of projections in Lemma 2.1. Then, the following equalities hold:
  1. (i)
    $p_{n}f_{n}q=qf_{n}p_{n}=0$, for $n\in\operatorname{\mathbb{Z}}$.
  2. (ii)
    $p_{n}f_{n\wedge k}p_{k}=p_{k}f_{n\wedge k}p_{n}=0$, for $n$, $k\in\operatorname{\mathbb{Z}}$, $n\neq k$, where $n\wedge k:=\min\left\{n,k\right\}$.
2.2. Noncommutative Calderón-Zygmund decomposition
For fixed $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$, we define the projection $\zeta_{s}$ as
(2.3) $\zeta_{s}=\bigg(\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}\bigg)^{\perp}.$
We require the following property of this projection, whose proof can be found in [2].
Lemma 2.3.
For fixed $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$, we have
$\phi({\textbf{1}}_{\operatorname{\mathcal{N}}}-\zeta_{s})\leq\frac{(2s+1)^{d}}{\lambda}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Below, we construct a slightly new noncommutative Calderón-Zygmund decomposition inspired by [2], and present its properties later.
Theorem 2.4.
Fix $f\in\operatorname{\mathcal{N}}_{c,+}$, $\lambda>0$, and $s\in\operatorname{\mathbb{N}}$. Let $(q_{n})_{n\in\operatorname{\mathbb{Z}}}$ and $(p_{n})_{n\in\operatorname{\mathbb{Z}}}$ be the sequences of projections in Lemma 2.1, where $q_{n}=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}q_{Q}\chi_{Q}$, $p_{n}=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}p_{Q}\chi_{Q}$. Suppose $h\in C_{c}^{\infty}(\operatorname{\mathbb{R}}^{d})$ satisfies
$\operatorname{supp}h\subset B_{1},\quad\int_{\operatorname{\mathbb{R}}^{d}}h=1,\quad h\geq 0.$
Define $h_{r}(x)=\frac{1}{r^{d}}h(\frac{x}{r})$. Then, $f$ admits the decomposition:
$f=g+b=g+\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0},$
where $g$ is positive and $b$ is self-adjoint in $\operatorname{\mathcal{N}}$. More precisely, these components are defined as follows:
$g=qfq+\sum\limits_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}p_{n},$
$\widetilde{b}_{\text{d}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{d},n}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{d},n}^{Q},\quad b_{\text{d}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{d},n}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{d},n}^{Q}-\widetilde{b}_{\text{d},n}^{Q}),$
$\widetilde{b}_{\text{off}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{off},n}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{off},n}^{Q},\quad b_{\text{off}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{off},n}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q}),$
where for each $Q\in\operatorname{\mathcal{Q}}_{n}$,
$b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\quad b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},$
$\widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)},\quad\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}.$
Proof.
[…]
$f=\sum_{n\in\widehat{\operatorname{\mathbb{Z}}}}\sum_{k\in\widehat{\operatorname{\mathbb{Z}}}}p_{n}f_{n\wedge k}p_{k}+\sum_{n\in\widehat{\operatorname{\mathbb{Z}}}}\sum_{k\in\widehat{\operatorname{\mathbb{Z}}}}p_{n}(f-f_{n\wedge k})p_{k}=:g+b.$
For the good function $g$, Lemma 2.2 implies
$g=qfq+\sum_{k\in\operatorname{\mathbb{Z}}}qf_{k}p_{k}+\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}q+\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k\in\operatorname{\mathbb{Z}}}p_{n}f_{n\wedge k}p_{k}=qfq+\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}f_{n}p_{n}.$
For the bad function $b$, we consider those terms with finite indices $n$ and $k$ satisfying $n\neq k$. Since $p_{k}=q_{k-1}-q_{k}$, a direct calculation shows
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k>n}p_{n}(f-f_{n})p_{k}=\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}(f-f_{n})(q_{n}-q),$
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{k<n}p_{n}(f-f_{k})p_{k}=\sum_{k\in\operatorname{\mathbb{Z}}}\sum_{n>k}p_{n}(f-f_{k})p_{k}=\sum_{k\in\operatorname{\mathbb{Z}}}(q_{k}-q)(f-f_{k})p_{k},$
which implies that
$b=\sum_{n\in\operatorname{\mathbb{Z}}}p_{n}(f-f_{n})p_{n}+\sum_{n\in\operatorname{\mathbb{Z}}}\big(p_{n}(f-f_{n})q_{n}+q_{n}(f-f_{n})p_{n}\big)$
by using Lemma 2.2 again. Next, we insert a convolution operator to decompose $b$ further. Let $h_{r}$ be defined as above. For a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, let
$\left\{\begin{array}[]{ll}b_{\text{d},n}^{Q}=p_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{d},n}^{Q}=b_{\text{d},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)},\end{array}\right.\quad\text{and}\quad\left\{\begin{array}[]{ll}b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q},\\ \widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}.\end{array}\right$.
By these definitions as well as the disjointness of dyadic cubes, we can express the bad function $b$ as
$\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{d},n}^{Q}}_{\widetilde{b}_{\text{d}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{d},n}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{d},n}^{Q}-\widetilde{b}_{\text{d},n}^{Q})}_{b_{\text{d}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{d},n}^{0}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\widetilde{b}_{\text{off},n}^{Q}}_{\widetilde{b}_{\text{off}}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}\widetilde{b}_{\text{off},n}}+\underbrace{\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q})}_{b_{\text{off}}^{0}=\sum\limits_{n\in\operatorname{\mathbb{Z}}}b_{\text{off},n}^{0}}.$
This yields the desired decomposition $f=g+b=g+\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0}$; the positivity of $g$ and self-adjointness of $b$ follows immediatly from their constructions. ∎
The decomposition in Theorem 2.4 admits the following properties:
Lemma 2.5.
Let $\zeta_{s}$ be the projection given as in (2.3). Then,
  1. (i)
    $\left\|g\right\|_{L_{1}(\operatorname{\mathcal{N}})}\leq\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$ and $\left\|g\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq 2^{d}\lambda$.
  2. (ii)
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}=0$, and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$. Furthermore, $\sum\limits_{n\in\operatorname{\mathbb{Z}}}\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
  3. (iii)
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{d},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
  4. (iv)
    For each $Q\in\operatorname{\mathcal{Q}}_{n}$, $\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{off},n}^{Q}=0$, and for any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)\widetilde{b}_{\text{off},n}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
  5. (v)
    For any $x,$ $y\in\operatorname{\mathbb{R}}^{d}$, $\zeta_{s}(x)b_{\text{off},n}^{0}(y)\zeta_{s}(x)=0$ when $|x-y|\leq(s-\sqrt{d})2^{-n}$.
Proof.
$(i)$ follows directly from Parcet’s noncommutative Calderón-Zygmund decomposition, as detailed in [1] or [2].
Next, consider $(ii)$. The proof contains three parts: the cancellation condition, the support condition, and the boundedness.
By the definition of $b_{\text{d},n}^{Q}$ in Theorem 2.4, we have $\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}=0$ for each $Q\in\operatorname{\mathcal{Q}}_{n}$. Since $\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}=1$, it follows that
$\int_{\operatorname{\mathbb{R}}^{d}}\widetilde{b}_{\text{d},n}^{Q}(x){\rm d}x=\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n}^{Q}(y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}(x-y){\rm d}x\Big){\rm d}y=0,$
which is the required cancellation property for $\widetilde{b}_{\text{d},n}^{Q}$.
Now we consider the support condition. For each $Q\in\operatorname{\mathcal{Q}}$, the construction of $\zeta_{s}$ in (2.3) implies
(2.4) $\zeta_{s}(x)p_{Q}=p_{Q}\zeta_{s}(x)=0\quad\text{if}\quad x\in(2s+1)Q.$
Indeed, if $x\in(2s+1)Q$, then
$\zeta_{s}(x)=1-\bigvee_{Q\in\operatorname{\mathcal{Q}}}p_{Q}\chi_{(2s+1)Q}(x)\leq 1-p_{Q},$
implying
$\zeta_{s}(x)p_{Q}=\zeta_{s}(x)(1-p_{Q})p_{Q}=0,$
and the same holds for $p_{Q}\zeta_{s}(x)$, which lead to (2.4). Then, note that $\operatorname{supp}(h_{\frac{\sqrt{d}}{2}\ell(Q)})\subset B_{\frac{\sqrt{d}}{2}\ell(Q)}$, we write $\zeta_{s}(x)\widetilde{b}_{\text{d},n}(y)\zeta_{s}(x)$ as follows:
$\zeta_{s}(x)\bigg(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}}b_{\text{d},n}^{Q}(u)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}u\bigg)\zeta_{s}(x)$
$\quad=\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\int_{u\in\big(B_{\frac{\sqrt{d}}{2}\ell(Q)}(y)\bigcap Q\big)}\Big(\zeta_{s}(x)p_{Q}(f-f_{n})(u)p_{Q}\zeta_{s}(x)\Big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}u.$
As mentioned, we let $Q_{u,n}$ be the unique cube in $\operatorname{\mathcal{Q}}_{n}$ that contains $u$. Hence, with (2.4), it is enough to show that
(2.5) $x\in(2s+1)Q_{u,n}\quad\text{if}\quad|x-y|\leq(s-\sqrt{d})2^{-n}.$
In fact, for $u\in B_{\frac{\sqrt{d}}{2}\ell(Q_{u,n})}(y)\bigcap Q_{u,n}$,
$|x-c(Q_{u,n})|$ $\leq|x-y|+|y-u|+|u-c(Q_{u,n})|$
$\leq(s-\sqrt{d})\ell(Q_{u,n})+\frac{\sqrt{d}}{2}\ell(Q_{u,n})+\frac{\sqrt{d}}{2}\ell(Q_{u,n})$
$=s\ell(Q_{u,n})$
implying (2.5).
The boundedness of $b_{\text{d},n}^{Q}$ results from the fact that $\phi(fp_{n})=\phi(f_{n}p_{n})$ together with (2.2). This means,
$\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\|b_{\text{d},n}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\leq\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\big(\phi(fp_{Q}\chi_{Q})+\phi(f_{n}p_{Q}\chi_{Q})\big)$
$=\sum_{n\in\operatorname{\mathbb{Z}}}\big(\phi(fp_{n})+\phi(f_{n}p_{n})\big)$
$\leq 2\phi(f).$
This completes the proof of $(ii)$.
$(iii)$ By the support condition in $(ii)$, we may restrict our attention to the nonconvolution part of $b_{\text{d},n}^{0}$. That is, by writing
$\zeta_{s}(x)b_{\text{d},n}^{0}\zeta_{s}(x)=\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\zeta_{s}(x)b_{\text{d},n}^{Q}(y)\zeta_{s}(x)-\sum\limits_{Q\in\operatorname{\mathcal{Q}}_{n}}\zeta_{s}(x)\widetilde{b}_{\text{d},n}^{Q}(y)\zeta_{s}(x),$
we only have to estimate the first sum, since the second one has already been considered in $(ii)$. The required estimate is obtained by using the same argument as in $(ii)$, and we therefore omit the details.
Finally, the proofs of $(iv)$ and $(v)$ follow from analogous reasoning and are also omitted.
Remark 2.6.
We will fix $s=[50\sqrt{d}]$ throughout this paper. Furthermore, the projection $\zeta_{s}$ is denoted simply by $\zeta$ when $s$ is fixed.
Remark 2.7.
Recall the construction of $q_{n}$: there exists an integer $n_{\lambda}(f)$ such that $q_{n}=1_{\operatorname{\mathcal{N}}}$ when $n\leq n_{\lambda}(f)$. Without loss of generality, we may set $n_{\lambda}(f)=0$. This leads to
$\widetilde{b}_{\text{d}}=\sum_{n\geq 1}\widetilde{b}_{\text{d},n},\quad\ b_{\text{d}}^{0}=\sum_{n\geq 1}b_{\text{d},n}^{0},\quad\ \widetilde{b}_{\text{off}}=\sum_{n\geq 1}\widetilde{b}_{\text{off},n},\quad\ b_{\text{off}}^{0}=\sum_{n\geq 1}b_{\text{off},n}^{0},$
thereby simplifying the summation.
3. Two reductions
This section introduces two key lemmas (Lemma 3.1 and Lemma 3.2), which reduce the proof of the main theorem to the cases of real kernels and lacunary sequences. For their proofs, we refer the reader to [7].
Step 1. Let $T_{\epsilon}$ be a noncommutative Calderón-Zygmund truncated operator with a complex kernel $K$ (defined as in (1.8)). We decompose $K$ into its real and imaginary parts: $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$. The truncated operators associated with $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$ are then given by
$\operatorname{Re\,}(T_{\epsilon})f(x)=\int_{|x-y|>\epsilon}\operatorname{Re\,}(K)(x,y)f(y){\rm d}y$
and
$\operatorname{Im\,}(T_{\epsilon})f(x)=\int_{|x-y|>\epsilon}\operatorname{Im\,}(K)(x,y)f(y){\rm d}y.$
Lemma 3.1.
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.7)) whose kernel $K$ satisfies (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}$ be defined as in (1.8). Then,
  1. (i)
    Both $\operatorname{Re\,}(K)$ and $\operatorname{Im\,}(K)$ satisfy (1.1) and (1.6).
  2. (ii)
    If $(T_{\epsilon})_{\epsilon>0}$ is of strong type $(p_{0},p_{0})$ for some $p_{0}\in(1,\infty)$, then, both $(\operatorname{Re\,}(T_{\epsilon}))_{\epsilon>0}$ and $(\operatorname{Im\,}(T_{\epsilon}))_{\epsilon>0}$ are of strong type $(p_{0},p_{0})$.
  3. (iii)
    If $(\operatorname{Re\,}(T_{\epsilon}))_{\epsilon>0}$ and $(\operatorname{Im\,}(T_{\epsilon}))_{\epsilon>0}$ are of weak type $(1,1)$, so is $(T_{\epsilon})_{\epsilon>0}$.
By this lemma, we may assume that the kernel $K$ is real throughout this paper.
Step 2. Let $\Phi$ be a smooth radial nonnegative function on $\operatorname{\mathbb{R}}^{d}$ such that
$\operatorname{supp}\Phi\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:1/2\leq|x|\leq 2\big\}\quad\text{and}\quad\sum_{j\in\operatorname{\mathbb{Z}}}\Phi_{j}(x)=1,\ x\in\operatorname{\mathbb{R}}^{d}\setminus\{0\},$
where $\Phi_{j}(x):=\Phi(\frac{x}{2^{j}\sqrt{d}})$. Thus, the truncated singular integrals $T_{\epsilon}f$ can be written as
$T_{\epsilon}f=\sum_{j\in\operatorname{\mathbb{Z}}}\int_{|x-y|>\epsilon}K(x,y)\Phi_{j}(x-y)f(y){\rm d}y.$
From the conditions $|x-y|>\epsilon$ and $\operatorname{supp}(\Phi_{j})\subset\{x\in\mathbb{R}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\}$, we see that for $i_{\epsilon}=\big[\log_{2}(\frac{\epsilon}{2\sqrt{d}})\big]+1$, the following equality holds (when $\epsilon>0$ is sufficiently small):
$T_{\epsilon}f(x)=\sum_{j\geq i_{\epsilon}}\int_{\operatorname{\mathbb{R}}^{d}}K(x,y)\Phi_{j}(x-y)f(y){\rm d}y+\int_{|x-y|>\epsilon}K(x,y)\Phi_{i_{\epsilon}}(x-y)f(y){\rm d}y.$
It is convenient to introduce the following notation:
(3.1) $K_{j}(x,y)=K(x,y)\Phi_{j}(x-y),$
(3.2) $S_{i}f(x)=\sum_{j\geq i}\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)f(y){\rm d}y,$
(3.3) $T_{\epsilon}^{i_{\epsilon}}f(x)=\int_{|x-y|>\epsilon}K_{i_{\epsilon}}(x,y)f(y){\rm d}y.$
Hence,
$T_{\epsilon}f(x)=S_{i_{\epsilon}}f(x)+T_{\epsilon}^{i_{\epsilon}}f(x).$
Lemma 3.2.
Let $T$ be a noncommutative Calderón-Zygmund operator (defined as in (1.7)) whose kernel $K$ satisfies (1.1) and the $L_{1}$-integral condition (1.6). Let $T_{\epsilon}^{i_{\epsilon}}$ be defined as in (3.3). Then
  1. (i)
    $(T_{\epsilon}^{i_{\epsilon}})_{\epsilon>0}$ is of strong type $(p,p)$ for $1<p\leq\infty$.
  2. (ii)
    $(T_{\epsilon}^{i_{\epsilon}})_{\epsilon>0}$ is of weak type $(1,1)$.
The preceding two lemmas imply that Theorem 1.1 follows directly from the weak type (1,1) boundedness of $S_{i}$ given in (3.2). Thus, it is enough to prove
Theorem 3.3.
Let $T$ be a noncommutative Calderón-Zygmund operator defined as in (1.7) with the real kernel $K$ satisfying (1.1) and the $L_{1}$-integral condition (1.6). Let $S_{i}$ be defined as in (3.2). Then, the sequence of operators $(S_{i})_{i\in\operatorname{\mathbb{Z}}}$ is of weak type $(1,1)$, that is, for any $f\in L_{1}(\operatorname{\mathcal{N}})$,
$\left\|(S_{i}f)_{i\in\operatorname{\mathbb{Z}}}\right\|_{\Lambda_{1,\infty}(\operatorname{\mathcal{N}};\ell_{\infty})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
More precisely, for any $f\in L_{1}(\operatorname{\mathcal{N}})$ and $\lambda>0$, there exists a projection $e\in\operatorname{\mathcal{N}}$ such that
$\sup_{i\in\operatorname{\mathbb{Z}}}\left\|eS_{i}fe\right\|_{L_{\infty(\operatorname{\mathcal{N}})}}\leq\lambda\quad\text{and}\quad\phi(e^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
4. The proof of Theorem 3.3
Note that every operator in a von Neumann algebra can be decomposed into a linear combination of four positive elements. By this fact and the density of $\operatorname{\mathcal{N}}_{c,+}$ in $L_{1}(\operatorname{\mathcal{N}})_{+}$, we may assume $f\in\operatorname{\mathcal{N}}_{c,+}$ without loss of generality. Now, consider a fixed $f\in\operatorname{\mathcal{N}}_{c,+}$ and $\lambda>0$. We apply Theorem 2.4 to decompose $f=g+b$. Then, by the quasi-triangle inequality, it suffices to find projections $e_{1},e_{2}\in\operatorname{\mathcal{N}}$ such that
(4.1) $\sup_{i\in\operatorname{\mathbb{Z}}}\left\|e_{1}S_{i}ge_{1}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda\quad\text{and}\quad\phi(e_{1}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
(4.2) $\sup_{i\in\operatorname{\mathbb{Z}}}\left\|e_{2}S_{i}be_{2}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda\quad\text{and}\quad\phi(e_{2}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Indeed, by setting $e=e_{1}\wedge e_{2}$, we obtain
$\left\|eS_{i}fe\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\left\|eS_{i}ge\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}+\left\|eS_{i}be\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq 2\lambda$
and
$\phi(e^{\bot})\leq\phi(e_{1}^{\bot})+\phi(e_{2}^{\bot})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which gives Theorem 3.3.
4.1. Estimate for the good function (4.1)
For any $i\in\operatorname{\mathbb{Z}}$, it follows from the definitions of $S_{i}$ and $T_{\epsilon}^{i_{\epsilon}}$ ((3.2) and (3.3)) that there exists $\epsilon_{i}>0$ such that
$S_{i}=S_{\epsilon_{i}}-T_{\epsilon_{i}}^{i}.$
Under the assumptions of Theorem 1.1 and Lemma 3.2 $(i)$, both $(S_{\epsilon_{i}})_{i\in\mathbb{Z}}$ and $(T_{\epsilon_{i}}^{i})_{i\in\mathbb{Z}}$ are of strong type $(p_{0},p_{0})$. Therefore, from the above equality, $(S_{i})_{i\in\mathbb{Z}}$ is also of strong type $(p_{0},p_{0})$. In addition, since $g$ is positive (by Theorem 2.4) and the kernel $K$ is real-valued (as stated in Section 3), the operator $S_{i}g$ is self-adjoint. These two facts imply that for any $i\in\operatorname{\mathbb{Z}}$, we can find a positive element $a\in\operatorname{\mathcal{N}}$ such that for any $i\in\operatorname{\mathbb{Z}}$,
$a\leq S_{i}g\leq a\quad\text{and}\quad\left\|a\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}\lesssim\left\|g\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}.$
Taking $e_{1}=\chi_{(0,\lambda]}(a)$. Then we have
$\lambda\leq-e_{1}ae_{1}\leq e_{1}S_{i}ge_{1}\leq e_{1}ae_{1}\leq\lambda.$
Finally, applying the Chebyshev inequality, the Hölder inequality, along with the boundedness of $g$ in Lemma 2.5 $(i)$, we obtain
$\phi(e_{1}^{\bot})\leq\lambda^{-p_{0}}\left\|a\right\|_{L_{p_{0}}(\operatorname{\mathcal{N}})}^{p_{0}}\leq\lambda^{-p_{0}}\left\|g\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}^{p_{0}-1}\left\|g\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which yields (4.1).
4.2. Estimate for the bad function (4.2)
We now give an estimate for the bad function $b$. First, we decompose $S_{i}b$ as
$S_{i}b=\zeta^{\bot}S_{i}b\zeta^{\bot}+\zeta S_{i}b\zeta^{\bot}+\zeta^{\bot}S_{i}b\zeta+\zeta S_{i}b\zeta,$
where the projection $\zeta$ is defined as in (2.3), and we omit its subscript as stated in Remark 2.6.
For each $j\in\operatorname{\mathbb{Z}}$, define the operator $\mathcal{T}_{j}$ by
$\mathcal{T}_{j}f(x)=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)f(y){\rm d}y.$
Then, $S_{i}$ admits a representation $S_{i}=\sum\limits_{j\geq i}\mathcal{T}_{j}$. The key step is to establish the following $L_{1}$-estimate:
(4.3) $\sum_{j\in\operatorname{\mathbb{Z}}}\left\|\zeta\mathcal{T}_{j}b\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
We claim that (4.3) holds, and use it to deduce (4.2). The proof of (4.3) itself will be given later.
Define the projection $\eta=\chi_{(0,\lambda]}\Big(\sum\limits_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\Big)$. It then follows that
$-\lambda\leq-\eta\sum_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\eta\leq\eta\zeta S_{i}b\zeta\eta\leq\eta\sum_{j\in\operatorname{\mathbb{Z}}}|\zeta\mathcal{T}_{j}b\zeta|\eta\leq\lambda.$
Moreover, by the Chebyshev inequality and (4.3),
$\phi(\eta^{\perp})\leq\lambda^{-1}\sum_{j\in\operatorname{\mathbb{Z}}}\left\|\zeta\mathcal{T}_{j}b\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Now, let $e_{2}=\zeta\wedge\eta$. Combining these results with Lemma 2.3, we obtain
$\left\|e_{2}S_{i}be_{2}\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\left\|\eta\zeta S_{i}b\zeta\eta\right\|_{L_{\infty}(\operatorname{\mathcal{N}})}\leq\lambda$
and
$\phi(e_{2})\leq\phi(\zeta^{\perp})+\phi(\eta^{\perp})\lesssim\lambda^{-1}\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which leads to (4.2).
It remains to prove (4.3). Recall from Theorem 2.4 that the bad function $b$ has the decomposition:
$b=\widetilde{b}_{\text{d}}+b_{\text{d}}^{0}+\widetilde{b}_{\text{off}}+b_{\text{off}}^{0}.$
Therefore, by the triangle inequality, it suffices to verify the following two lemmas:
Lemma 4.1.
$(i)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$; $(ii)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}b^{0}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
Lemma 4.2.
$(i)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$; $(ii)$ $\sum\limits_{j\in\operatorname{\mathbb{Z}}}\big\|\zeta\mathcal{T}_{j}b_{\text{off}}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}$.
We first introduce a lemma that will be essential for proving Lemma 4.1 and Lemma 4.2. For a fixed dyadic cube $Q$, we recall the kernel average
$\mathcal{K}_{Q}(x,y)=\frac{1}{|Q|}\int_{|v|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+v)-K(x,y)|{\rm d}v.$
When there is no risk of confusion, it is convenient to denote by $\mathcal{K}_{j,Q}$ the corresponding average for the truncated kernel $K_{j}$:
$\mathcal{K}_{j,Q}(x,y):=\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z.$
Lemma 4.3.
Fix a dyadic cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$ and $j\in\operatorname{\mathbb{Z}}$. Then, for $n\geq 3$,
$\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y){\rm d}x\lesssim\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1)+2^{-n}.$
Proof.
From the definition of $K_{j}$ in (3.1), we have
$\mathcal{K}_{j,Q}(x,y)$ $\leq\frac{1}{|Q|}\Big(\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)-K(x,y)||\Phi_{j}(x-y)|{\rm d}z$
$\quad+\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)||\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|{\rm d}z\Big).$
Consequently,
$\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y){\rm d}x\leq A_{1}+A_{2},$
where
$A_{1}$ $:=\frac{1}{|Q|}\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\Big(\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)-K(x,y)|{\rm d}z|\Phi_{j}(x-y)|\Big){\rm d}x$
$=\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in Q}\Big(\mathcal{K}_{Q}(x,y)|\Phi_{j}(x-y)|\Big){\rm d}x,$
and
$A_{2}:=\frac{1}{|Q|}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\sup_{y\in Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K(x,y+z)||\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)|{\rm d}z\Big){\rm d}x.$
We first estimate $A_{1}$. When $y\in Q$, by $\operatorname{supp}(\Phi_{j})\subset\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$ and
(4.4) $\frac{\sqrt{d}}{2}\ell(Q)=2^{-n+j-1}\sqrt{d}\leq 2^{j-4}\sqrt{d},\quad\text{when}\ n\geq 3,$
we obtain
$|x-c(Q)|\leq|x-y|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+(\sqrt{d}/2)\ell(Q)\leq 2^{j+2}\sqrt{d},$
$|x-c(Q)|\geq|x-y|-|y-c(Q)|\leq 2^{j-1}\sqrt{d}-(\sqrt{d}/2)\ell(Q)\leq 2^{j-2}\sqrt{d}.$
This means,
$2^{n-2}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+2}\sqrt{d}\ell(Q).$
Hence,
$A_{1}$ $\leq\int_{2^{n-2}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+2}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}(x,y){\rm d}x$
$=\sum_{k=-2}^{1}\int_{2^{n+k}\sqrt{d}\ell(Q)\leq|x-c(Q)|\leq 2^{n+k+1}\sqrt{d}\ell(Q)}\sup_{y\in Q}\mathcal{K}(x,y){\rm d}x$
$\leq\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1).$
Consider $A_{2}$. We first claim that there exists a constant $C_{d}$ such that the function $x\mapsto\big(\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\big)$ is supported in the set
(4.5) $\big\{x\in\operatorname{\mathbb{R}}^{d}:|x-c(Q)|\leq C_{d}2^{j}\big\}=:E_{j,Q},$
whenever $y\in Q$ and $z\in B_{\frac{\sqrt{d}}{2}\ell(Q)}$. To prove this, by the support of $\Phi_{j}$, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, at least one of the following conditions holds:
  1. $(i)$
    $2^{j-1}\sqrt{d}\leq|x-(y+z)|\leq 2^{j+1}\sqrt{d}$,
  2. $(ii)$
    $2^{j-1}\sqrt{d}\leq|x-y|\leq 2^{j+1}\sqrt{d}$.
In case $(i)$, it follows from $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)$, $y\in Q$ and (4.4) that
$|x-c(Q)|\leq|x-(y+z)|+|z|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d}+2^{j-4}\sqrt{d}.$
In case $(ii)$, similarly,
$|x-c(Q)|\leq|x-y|+|y-c(Q)|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d}.$
Thus (4.5) holds.
Moreover, when $\Phi_{j}(x-(y+z))-\Phi_{j}(x-y)\neq 0$, we have another useful estimate:
(4.6) $|x-(y+z)|\approx 2^{j}.$
Indeed, this is immediate in case $(i)$. And in case $(ii)$, using $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)$ and (4.4) again, we have
$|x-(y+z)|\leq|x-y|+|z|\leq 2^{j+1}\sqrt{d}+2^{j-4}\sqrt{d},$
$|x-(y+z)|\geq|x-y|-|z|\geq 2^{j-1}\sqrt{d}-2^{j-4}\sqrt{d},$
which also implies $|x-(y+z)|\approx 2^{j}$, leading to (4.6).
Now, we can bound $A_{2}$. By the size condition (1.1), the mean-value theorem, the bound $|z|\leq\frac{\sqrt{d}}{2}\ell(Q)=2^{-n+j-1}\sqrt{d}$, together with (4.5) and (4.6),
$A_{2}$ $\lesssim\int_{E_{j,Q}}\Big(\sup_{y\in Q}\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\frac{|z|2^{-j}}{|x-(y+z)|^{d}}{\rm d}z\Big){\rm d}x$
$\lesssim 2^{-n-jd}\int_{E_{j,Q}}\Big(\sup_{y\in Q}\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}1{\rm d}z\Big){\rm d}x$
$\lesssim 2^{-n}.$
This completes the proof. ∎
4.2.1. The proof of Lemma 4.1
$(i)$ We first observe that, for any $j\in\operatorname{\mathbb{Z}}$,
(4.7) $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta=0,\quad\text{when}\ n\leq 3.$
Indeed, for a fixed $x\in\operatorname{\mathbb{R}}^{d}$, Lemma 2.5 $(ii)$ and Remark 2.6 ($s=[50\sqrt{d}]$) implies
$\zeta(x)\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)\zeta(x)$
$\quad=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\zeta(x)\widetilde{b}_{\text{d},n-j}(y)\zeta(x)\Big)\chi_{\big\{|x-y|>([50\sqrt{d}]-\sqrt{d})2^{-n+j}\big\}}(y){\rm d}y.$
Since $\operatorname{supp}(\Phi_{j})\subseteq\big\{x\in\operatorname{\mathbb{R}}^{d}:2^{j-1}\sqrt{d}\leq|x|\leq 2^{j+1}\sqrt{d}\big\}$, we get
$K_{j}(x,y)=K(x,y)\Phi_{j}(x-y)=0,$
when
$|x-y|\geq([50\sqrt{d}]-\sqrt{d})2^{-n+j}\geq([50\sqrt{d}]-\sqrt{d})2^{j-3}\geq 2^{j+1}\sqrt{d},$
where the second inequality uses $n\leq 3$. This shows that $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta=0$ when $n\leq 3$. The same argument works for $\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta$, giving rise to (4.7).
We begin with estimating $\widetilde{b}_{\text{d}}$. By Remark 2.7 and (4.7), it suffices to establish
$\sum_{n\geq 4}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
It follows from the definition of $\widetilde{b}_{\text{d},n}$ in Theorem 2.4, the cancellation condition in Lemma 2.5 $(ii)$, and the definition of $\widetilde{b}_{\text{d},{n-j}}^{Q}$ in Theorem 2.4 that
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}(x)$ $=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)\widetilde{b}_{\text{d},{n-j}}^{Q}(y){\rm d}y$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q)\big)b_{\text{d},n-j}^{Q}(u)h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}y{\rm d}u.$
Then, by $\operatorname{supp}(h_{\frac{\sqrt{d}}{2}\ell(Q)})\subseteq B_{\frac{\sqrt{d}}{2}\ell(Q)}$, $\operatorname{supp}(b^{Q}_{\text{d},n-j})\subseteq Q$, and using the Minkowski inequality,
$\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{\left\|h\right\|_{\infty}}{\ell(Q)^{d}}\int_{Q}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y{\rm d}u\Big){\rm d}x$
$\lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\Big(\int_{\operatorname{\mathbb{R}}^{d}}\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y{\rm d}x\Big)\|b_{\text{d},n-j}^{Q}(u)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u.$
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. For $y\in B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)$, we have
$|y-c(Q)|\leq|y-u|+|u-c(Q)|\leq\frac{\sqrt{d}}{2}\ell(Q)+\frac{\sqrt{d}}{2}\ell(Q)\leq\sqrt{d}\ell(Q)$
whenever $u\in Q$. Hence, by letting z=yc(Q)z=y-c(Q),
$\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y$
(4.8) $\quad\lesssim\frac{1}{|Q|}\int_{|z|\leq\sqrt{d}\ell(Q)}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}z.$
Now, let the cube $\widehat{Q}$ be the unique father cube of $Q$ in $\operatorname{\mathcal{Q}}_{n-j-1}$. This implies
[…]
$\leq\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y),$
which leads to
$\int_{\operatorname{\mathbb{R}}^{d}}\frac{1}{|Q|}\int_{B_{\frac{\sqrt{d}}{2}\ell(Q)}(u)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y{\rm d}x$
$\quad\lesssim\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y){\rm d}x$
$\quad\lesssim\mathcal{H}(n-3)+\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1},$
where the last bound comes from Lemma 4.3 with $n$ replaced by n1n-1. Finally, applying the $L_{1}$-integral condition (1.6) and the boundedness of $b_{\text{d},n}^{Q}$ in Lemma 2.5 $(ii)$, we obtain
$\sum_{n\geq 4}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{d}}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\lesssim\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\big\|b_{\text{d},n-j}^{Q}(u)\big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}u$
$\quad\quad\left(\mathcal{H}(n-3)+\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1}\right)$
$\ \lesssim\sum_{n\geq 4}\big(4\mathcal{H}(n-3)+2^{-n+1}\big)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\big\|b_{\text{d},m}^{Q}\big\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\ \lesssim\left\|f\right\|_{1}.$
$(ii)$ Again by Remark 2.7 and (4.7), it suffices to prove:
$\sum_{n\geq 4}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. From the definition of $\widetilde{b}_{\text{d},n-j}^{Q}$ in Theorem 2.4,
$\mathcal{T}_{j}\widetilde{b}_{\text{d},n-j}^{Q}(x)$ $=\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)\Big(\int_{\operatorname{\mathbb{R}}^{d}}b_{\text{d},n-j}^{Q}(y-z)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big){\rm d}y$
$=\int_{\operatorname{\mathbb{R}}^{d}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y+z)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big)b_{\text{d},n-j}^{Q}(y){\rm d}y.$
Moreover, by $\int_{\operatorname{\mathbb{R}}^{d}}h_{\frac{\sqrt{d}}{2}\ell(Q)}=1$,
$\mathcal{T}_{j}b_{\text{d},n-j}^{Q}(x)=\int_{\operatorname{\mathbb{R}}^{d}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z\Big)b_{\text{d},n-j}^{Q}(y){\rm d}y.$
Noting again the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$, $b_{\text{d},n-j}^{Q}$ and using the Minkowski inequality, we get
$\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\ \lesssim\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{\operatorname{\mathbb{R}}^{d}}\Big(\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\Big){\rm d}x\|b_{\text{d},n-j}^{Q}(y)\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}y$
$\ \lesssim\big(\mathcal{H}(n-2)+\mathcal{H}(n-1)+\mathcal{H}(n)+\mathcal{H}(n+1)+2^{-n}\big)\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\|b_{\text{d},n-j}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})},$
where the last inequality follows from Lemma 4.3. Finally, an application of the $L_{1}$-integral condition (1.6), and the boundedness of $b_{\mathrm{d},n}^{Q}$ in Lemma 2.5 $(ii)$, yield
$\sum_{n\geq 4}\sum_{j\leq n-1}\left\|\zeta\mathcal{T}_{j}b_{\text{d},n-j}^{0}\zeta\right\|_{L_{1}(\operatorname{\mathcal{N}})}$ $\lesssim\sum_{n\geq 4}\left(4\mathcal{H}(n-2)+2^{-n}\right)\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\|b_{\text{d},m}^{Q}\|_{L_{1}(\operatorname{\mathcal{N}})}$
$\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})},$
which completes the argument.
4.2.2. The proof of Lemma 4.2
$(i)$ As in the proof for the diagonal part $b_{\text{d}}$, we state the following equality, whose proof is similar to that of (4.7): for any $j\in\operatorname{\mathbb{Z}}$,
(4.9) $\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta=\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta=0,\quad\text{when}\ n\leq 3.$
Together with Remark 2.7, it suffices to prove
$\sum_{n\geq 4}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
For a fixed $x\in\operatorname{\mathbb{R}}^{d}$, the definition of $\widetilde{b}_{\text{off},n-j}$ in Theorem 2.4 and the cancellation condition in Lemma 2.5 $(iv)$ give
$\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}(x)=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q}(y){\rm d}y.$
Next, recall the definition of $\widetilde{b}_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $\widetilde{b}_{\text{off},n}^{Q}=b_{\text{off},n}^{Q}*h_{\frac{\sqrt{d}}{2}\ell(Q)}$, where $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Accordingly, we can majorize $\sum\limits_{n\geq 4}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}\widetilde{b}_{\text{off},n-j}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ by the sum of the following four terms:
$F_{i}:=\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{\operatorname{\mathbb{R}}^{d}}\big(K_{j}(x,y)-K_{j}(x,c(Q))\big)\widetilde{b}_{\text{off},n-j}^{Q,i}(y){\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, where
$\widetilde{b}_{\text{off},n-j}^{Q,1}:=p_{Q}\big((f\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)q_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,2}:=p_{Q}\big((f_{Q}\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)q_{Q},$
$\widetilde{b}_{\text{off},n-j}^{Q,3}:=q_{Q}\big((f\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)p_{Q},\quad\widetilde{b}_{\text{off},n-j}^{Q,4}:=q_{Q}\big((f_{Q}\chi_{Q})*h_{\frac{\sqrt{d}}{2}\ell(Q)}\big)p_{Q}.$
We provide a detailed proof only for $F_{1}$; analogous arguments apply to $F_{2}$$F_{4}$ under the fact that $\phi(f_{n}p_{n})=\phi(fp_{n})$.
Define
$\mathcal{L}_{j}(x,u):=\int_{|u-y|\leq\frac{\sqrt{d}}{2}\ell(Q)}(K_{j}(x,y)-K_{j}(x,c(Q)))h_{\frac{\sqrt{d}}{2}\ell(Q)}(y-u){\rm d}y,$
and decompose it into positive and negative parts:
$\mathcal{L}_{j}(x,u)=\mathcal{L}_{j}^{+}(x,u)-\mathcal{L}_{j}^{-}(x,u).$
Hence, for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$, using the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm inside $F_{1}$ equals
$\Big\|\int_{Q}\mathcal{L}_{j}(x,u)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
By the triangle inequality, it is enough to treat the positive part, i.e.
(4.10) $\Big\|\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)q_{Q}{\rm d}u\Big\|_{L_{1}(\operatorname{\mathcal{M}})}.$
Let $\widehat{Q}$ be the unique father cube of $Q$ in $\operatorname{\mathcal{Q}}_{n-j-1}$. As in the proof of Lemma 4.1 $(i)$, when $u\in Q$, we can estimate $\mathcal{L}_{j}^{+}(x,u)$ as follows:
$|\mathcal{L}_{j}^{+}(x,u)|$ $\lesssim\frac{1}{|Q|}\int_{|u-y|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y)-K_{j}(x,c(Q))|{\rm d}y$
$\lesssim\frac{1}{|\widehat{Q}|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(\widehat{Q})}|K_{j}(x,z+c(Q))-K_{j}(x,c(Q))|{\rm d}z$
(4.11) $\leq\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y).$
By Lemma A.4 in Appendix A.3, there exists a contraction $m(x)$ such that
$\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)q_{Q}{\rm d}u$
$\quad=\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}m(x)\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)q_{Q}f(u)q_{Q}d{\rm d}u\Big)^{\frac{1}{2}}$
$\quad=\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}p_{Q}m(x)\Big(\int_{Q}\mathcal{L}^{+}_{j}(x,u)q_{Q}f(u)q_{Q}d{\rm d}u\Big)^{\frac{1}{2}}.$
Here we used that $(p_{Q}Xp_{Q})^{\frac{1}{2}}=(p_{Q}Xp_{Q})^{\frac{1}{2}}p_{Q}$ when $X$ is positive. Together it with the Hölder inequality and (4.2.2), we can bound (4.10) by
$\Big\|\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}\Big\|p_{Q}m(x)\Big(\int_{Q}\mathcal{L}_{j}^{+}(x,u)q_{Q}f(u)q_{Q}{\rm d}u\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}$
$\lesssim\Big(\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y)\Big)\tau\Big(\int_{Q}p_{Q}f(u)p_{Q}{\rm d}u\Big)^{\frac{1}{2}}\tau\Big(p_{Q}m(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)m(x)^{*}p_{Q}\Big)^{\frac{1}{2}}.$
On the one hand, it is easy to see
$\tau\Big(\int_{Q}p_{Q}f(u)p_{Q}{\rm d}u\Big)\leq\phi\big(fp_{Q}\chi_{Q}\big).$
On the other hand, by the definition of $f_{Q}$, (2.1) and $|m(x)|^{2}\leq 1$,
$\tau\Big(p_{Q}m(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)m(x)^{*}p_{Q}\Big)$ $=\tau\big(|Q|p_{Q}m(x)q_{Q}f_{Q}q_{Q}m(x)^{*}p_{Q}\big)$
$\leq\lambda|Q|\tau\big(p_{Q}m(x)q_{Q}m(x)^{*}p_{Q}\big)$
$\leq\lambda|Q|\tau(p_{Q}).$
Hence,
[…]
This estimate, combined with Lemma 4.3 and the $L_{1}$-integral condition (1.6), implies
$F_{1}$ $\lesssim\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\Big(\int_{\operatorname{\mathbb{R}}^{d}}\sup_{y\in\widehat{Q}}\mathcal{K}_{j,\widehat{Q}}(x,y){\rm d}x\Big)\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}$
$\lesssim\sum_{n\geq 4}\Big(\mathcal{H}(n-3)+\mathcal{H}(n-3)+\mathcal{H}(n-1)+\mathcal{H}(n)+2^{-n+1}\Big)$
$\quad\quad\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}$
$\lesssim\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}.$
Finally, we prove
(4.12) $\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}\lesssim\|f\|_{L_{1}(\operatorname{\mathcal{N}})},$
which gives the desired bound for $F_{1}$. Indeed, by applying the Cauchy-Schwarz inequality twice, we estimate the left-hand side of (4.12) as
$\sum_{n\in\operatorname{\mathbb{Z}}}\Big(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\phi(fp_{Q}\chi_{Q})\Big)^{\frac{1}{2}}\Big(\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\tau(p_{Q})(|Q|\lambda\Big)^{\frac{1}{2}}$
$\quad\leq\Big(\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\phi(fp_{Q}\chi_{Q})\Big)^{\frac{1}{2}}\Big(\sum_{n\in\operatorname{\mathbb{Z}}}\sum_{Q\in\operatorname{\mathcal{Q}}_{n}}\tau(p_{Q})(|Q|\lambda)\Big)^{\frac{1}{2}}$
$\quad=\Big(\sum_{n\geq 1}\phi(fp_{n})\Big)^{\frac{1}{2}}\Big(\sum_{n\geq 1}\lambda\phi(p_{n})\Big)^{\frac{1}{2}}.$
Then, it follows from (2.2) that $\sum\limits_{n\geq 1}\phi(p_{n})=\phi(1-q)$ and $\sum\limits_{n\geq 1}\phi(fp_{n})=\phi\big(f(1-q)\big)$. Combined with Lemma 2.1 $(iv)$, this yields
$\sum_{m\geq 1}\sum_{Q\in\operatorname{\mathcal{Q}}_{m}}\phi\big(fp_{Q}\chi_{Q}\big)^{\frac{1}{2}}\big(\lambda|Q|\tau(p_{Q})\big)^{\frac{1}{2}}\leq\Big(\lambda\phi(1-q)\Big)^{\frac{1}{2}}\Big(\phi(f(1-q))\Big)^{\frac{1}{2}}\lesssim\|f\|_{L_{1}(\operatorname{\mathcal{N}})},$
which proves (4.12) and completes the estimate for $F_{1}$.
$(ii)$ We now consider $b_{\text{off}}^{0}$. By (4.9) and Remark 2.7, it suffices to prove
$\sum_{n\geq 4}\sum_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}\lesssim\left\|f\right\|_{L_{1}(\operatorname{\mathcal{N}})}.$
For a fixed $x\in\operatorname{\mathbb{R}}^{d}$, from the definition of $b_{\text{off},n-j}^{0}$ in Theorem 2.4, a direct calculation implies that $\mathcal{T}_{j}b_{\text{off},n-j}^{0}(x)$ equals to
$\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{\operatorname{\mathbb{R}}^{d}}K_{j}(x,y)(b_{\text{off},n}^{Q}-\widetilde{b}_{\text{off},n}^{Q})(y){\rm d}y$
$=\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}\int_{Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y)-K_{j}(x,y+z)\big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)b^{Q}_{\text{off},n-j}(y){\rm d}z{\rm d}y.$
Recall the definition of $b_{\text{off},n}^{Q}$ in Theorem 2.4: for a fixed cube $Q\in\operatorname{\mathcal{Q}}_{n}$, $b_{\text{off},n}^{Q}=p_{Q}(f-f_{n})q_{Q}\chi_{Q}+q_{Q}(f-f_{n})p_{Q}\chi_{Q}$. Again, $\sum\limits_{n\geq 4}\sum\limits_{j\leq n-1}\big\|\zeta\mathcal{T}_{j}b_{\text{off},n-j}^{0}\zeta\big\|_{L_{1}(\operatorname{\mathcal{N}})}$ can be majorized by the sum of the following four terms:
$G_{i}:=$ $\sum_{n\geq 4}\sum_{j\leq n-1}\sum_{Q\in\operatorname{\mathcal{Q}}_{n-j}}$
$\int_{\operatorname{\mathbb{R}}^{d}}\Big\|\int_{Q}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)b_{\text{off},n-j}^{Q,i}(y,z){\rm d}z{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})}{\rm d}x,$
for $i=1,2,3,4$, where
$b_{\text{off},n-j}^{Q,1}(y,z)=p_{Q}(f\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)q_{Q},\quad b_{\text{off},n-j}^{Q,2}(y,z)=p_{Q}(f_{Q}\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)q_{Q},$
$b_{\text{off},n-j}^{Q,3}(y,z)=q_{Q}(f\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)p_{Q},\quad b_{\text{off},n-j}^{Q,4}(y,z)=q_{Q}(f_{Q}\chi_{Q})(y)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z)p_{Q}.$
We estimate only $G_{1}$; the other terms are similar by using $\phi(f_{n}p_{n})=\phi(fp_{n})$.
Fix a cube $Q\in\operatorname{\mathcal{Q}}_{n-j}$. By the supports of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ and $\chi_{Q}$, the noncommutative $L_{1}$-norm inside $G_{1}$ equals
$\Big\|\int_{Q}\mathcal{J}_{j}(x,y)p_{Q}f(y)q_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})},$
where
$\mathcal{J}_{j}(x,y):=\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}\big(K_{j}(x,y+z)-K_{j}(x,y)\big)h_{\frac{\sqrt{d}}{2}\ell(Q)}(z){\rm d}z.$
We also write $\mathcal{J}_{j}(x,y)=\mathcal{J}^{+}_{j}(x,y)-\mathcal{J}^{-}_{j}(x,y)$ and consider only $\mathcal{J}^{+}_{j}(x,y)$, i.e.
(4.13) $\Big\|\int_{Q}\mathcal{J}_{j}^{+}(x,y)p_{Q}f(y)q_{Q}{\rm d}y\Big\|_{L_{1}(\operatorname{\mathcal{M}})},$
Immediately, when $y\in Q$, it follows from the definition of $h_{\frac{\sqrt{d}}{2}\ell(Q)}$ that
$|\mathcal{J}^{+}_{j}(x,y)|\lesssim\frac{1}{|Q|}\int_{|z|\leq\frac{\sqrt{d}}{2}\ell(Q)}|K_{j}(x,y+z)-K_{j}(x,y)|{\rm d}z\leq\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y).$
Now, following the same reasoning as for $F_{1}$ in Lemma 4.2 $(i)$, we bound (4.13) by
$\Big\|\Big(\int_{Q}\mathcal{J}_{j}^{+}(x,y)p_{Q}f(y)p_{Q}{\rm d}y\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}\Big\|p_{Q}m^{\prime}(x)\Big(\int_{Q}\mathcal{J}_{j}^{+}(x,y)q_{Q}f(y)q_{Q}{\rm d}y\Big)^{\frac{1}{2}}\Big\|_{L_{2}(\operatorname{\mathcal{M}})}$
$\lesssim\Big(\sup_{y\in Q}\mathcal{K}_{j,Q}(x,y)\Big)\tau\Big(\int_{Q}p_{Q}f(y)p_{Q}{\rm d}y\Big)^{\frac{1}{2}}\tau\Big(p_{Q}m^{\prime}(x)\big(\int_{Q}q_{Q}f(u)q_{Q}{\rm d}u\big)(m^{\prime}(x))^{*}p_{Q}\Big)^{\frac{1}{2}},$
where $m^{\prime}(x)$ is the contraction provided by Lemma A.4 in Appendix A.3. Finally, the remaining part of the proof follows the same method as used for $F_{1}$. Hence, we complete the estimate for $G_{1}$ and the proof of Lemma 4.2.
Appendix A Noncommutative $L_{p}$ spaces
This part consists of three subsections: noncommutative $L_{p}$ spaces, noncommutative maximal norms, and noncommutative square functions. All these content are standard in noncommutative harmonic analysis, we refer to see [5, 9, 17, 18, 19] for more details.
A.1. Noncommutative $L_{p}$ spaces
[…]
Let $0<p<\infty$. For any $x\in\operatorname{\mathcal{S}}$, the operator $|x|^{p}\in\operatorname{\mathcal{S}}_{+}$ (meaning $\tau(|x|^{p})<\infty$) , where $|x|=(x^{*}x)^{\frac{1}{2}}$ is the modulus of $x$. We define
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}})}=\big(\tau(|x|^{p})\big)^{\frac{1}{p}},\quad x\in\operatorname{\mathcal{S}}.$
The noncommutative $L_{p}$ space associated with $(\operatorname{\mathcal{M}},\tau)$, denoted by $L_{p}(\operatorname{\mathcal{M}},\tau)$ or simply $L_{p}(\operatorname{\mathcal{M}})$, is defined as the completion of the space $(\operatorname{\mathcal{S}},\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}})})$. For convenience, we set $L_{\infty}(\operatorname{\mathcal{M}})=\operatorname{\mathcal{M}}$ equipped with the operator norm $\left\|\cdot\right\|_{\operatorname{\mathcal{M}}}$. We refer the reader to [19] for more information regarding noncommutative $L_{p}$ spaces.
A.2. Noncommutative maximal norms
The generalization of the maximal function to noncommutative spaces faces a challenge: operators in von Neumann algebras are not directly comparable. Remarkably, Pisier [18] and Junge [9] overcame this difficulty by directly defining the noncommutative maximal norm.
Definition A.1.
For $1\leq p\leq\infty$, we define $L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ as the space of all sequences $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}$ in $L_{p}(\operatorname{\mathcal{M}})$ which admit a factorization of the following form: there exist $a$, $b\in L_{2p}(\operatorname{\mathcal{M}})$ and a bounded sequence $y=(y_{n})_{n\in\operatorname{\mathbb{Z}}}$ in $L_{\infty}(\operatorname{\mathcal{M}})$ such that
$x_{n}=ay_{n}b,\quad n\in\operatorname{\mathbb{Z}}.$
The norm of $x\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ is defined as
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf_{x_{n}=ay_{n}b}\big\{\left\|a\right\|_{L_{2p}(\operatorname{\mathcal{M}})}\sup_{n\in\operatorname{\mathbb{Z}}}\left\|y_{n}\right\|_{L_{\infty}(\operatorname{\mathcal{M}})}\left\|b\right\|_{L_{2p}(\operatorname{\mathcal{M}})}\big\},$
where the infimum is taken over all factorizations of $x$ as above.
One can check that $L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ is a Banach space equipped with the noncommutative maximal norm $\left\|\cdot\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}$. A positive sequence $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ if and only if there exists a positive element $a\in L_{p}(\operatorname{\mathcal{M}})$ such that
$0<x_{n}\leq a,\quad\forall n\in\operatorname{\mathbb{Z}}.$
Moreover, its norm is given by
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf\big\{\left\|a\right\|_{L_{p}(\operatorname{\mathcal{M}})}:0<x_{n}\leq a,\ \forall n\in\operatorname{\mathbb{Z}}\big\}.$
Similarly, when $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}$ is a sequence of self-adjoint operators in $L_{p}(\operatorname{\mathcal{M}})$, $x\in L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})$ if and only if there exists a positive element $a\in L_{p}(\operatorname{\mathcal{M}})$ such that
$-a\leq x_{n}\leq a,\quad\forall n\in\operatorname{\mathbb{Z}}.$
In this case,
$\left\|x\right\|_{L_{p}(\operatorname{\mathcal{M}};\ell_{\infty})}=\inf\big\{\left\|a\right\|_{L_{p}(\operatorname{\mathcal{M}})}:-a\leq x_{n}\leq a,\ a>0,\ \forall n\in\operatorname{\mathbb{Z}}\big\}.$
The concept of the weak maximal norm is also essential in this paper.
Definition A.2.
For $1\leq p<\infty$ and $x=(x_{n})_{n\in\operatorname{\mathbb{Z}}}\subset L_{p}(\operatorname{\mathcal{M}})$, we define
$\left\|x\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\lambda\inf_{e\in\operatorname{\mathcal{P}}(\operatorname{\mathcal{M}})}\big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:\left\|ex_{n}e\right\|_{L_{\infty}(\operatorname{\mathcal{M}})}\leq\lambda,\ \forall n\in\operatorname{\mathbb{Z}}\big\},$
[…]
The space $\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})$ is a quasi-Banach space. For a positive sequence $x=(x_{n})\subset L_{p}(\operatorname{\mathcal{M}})$, the weak norm also admits the equivalent expression:
$\left\|x\right\|_{\Lambda_{p,\infty}(\operatorname{\mathcal{M}};\ell_{\infty})}=\sup_{\lambda>0}\lambda\inf_{e\in\operatorname{\mathcal{P}}(\operatorname{\mathcal{M}})}\Big\{\big(\tau(e^{\perp})\big)^{\frac{1}{p}}:0\leq ex_{n}e\leq\lambda,\ \forall n\in\operatorname{\mathbb{Z}}\Big\}.$
A.3. Noncommutative square functions
Consider a measure space $(\Omega,\mu)$ and an operator-valued function $f$. Define
$\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}=\Big\|\Big(\int_{\Omega}f^{*}(t)f(t){\rm d}\mu(t)\Big)^{\frac{1}{2}}\Big\|_{L_{p}(\operatorname{\mathcal{M}})}.$
Then, the column space $L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))$ is defined as the space of all such functions $f$ for which $\left\|f\right\|_{L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Omega))}<\infty$. When the measure is clear, we write simply ${\rm d}\omega$ instead of ${\rm d}\mu(\omega)$.
First, we introduce a Hölder inequality for the square functions.
Lemma A.3.
Let $0<p,q,r\leq\infty$ be such that $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$. Then for any $f\in L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$ and $g\in L_{q}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$,
$\Big\|\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega\Big\|_{L_{r}(\operatorname{\mathcal{M}})}\leq\Big\|\Big(\int_{\Sigma}|f(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}\Big\|_{L_{p}(\operatorname{\mathcal{M}})}\Big\|\Big(\int_{\Sigma}|g(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}\Big\|_{L_{q}({\mathcal{M}})}.$
The next result was communicated to the authors of [7] by Cadilhac during their work on the maximal Calderón-Zygmund operators. We include a proof for completeness.
Lemma A.4.
Let $0<p,q\leq\infty$. For any $f\in L_{p}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$ and $g\in L_{q}(\operatorname{\mathcal{M}};L_{2}^{c}(\Sigma))$, there exists a contraction operator $m$ such that
$\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=\Big(\int_{\Sigma}f^{*}(\omega)f(\omega){\rm d}\omega\Big)^{\frac{1}{2}}m\Big(\int_{\Sigma}g^{*}(\omega)g(\omega){\rm d}\omega\Big)^{\frac{1}{2}}.$
Proof.
Fix $\epsilon>0$ and set
$F_{\epsilon}:=\int_{\Sigma}f^{*}(\omega)f(\omega){\rm d}\omega+\epsilon{\textbf{1}}_{\operatorname{\mathcal{M}}}\quad\text{and}\quad G_{\epsilon}:=\int_{\Sigma}g^{*}(\omega)g(\omega){\rm d}\omega+\epsilon{\textbf{1}}_{\operatorname{\mathcal{M}}}.$
Then, we define $m_{\epsilon}$ by
$m_{\epsilon}:=F_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega\Big)G_{\epsilon}^{-\frac{1}{2}},$
so that
(A.1) $\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=F_{\epsilon}^{\frac{1}{2}}m_{\epsilon}G_{\epsilon}^{\frac{1}{2}}.$
[…]
$m_{\epsilon}=\int_{\Sigma}\big(f(\omega)F_{\epsilon}^{-\frac{1}{2}}\big)^{*}\big(g(\omega)G_{\epsilon}^{-\frac{1}{2}}\big){\rm d}\omega.$
Applying Lemma A.3, we get
$\|m_{\epsilon}\|_{L_{\infty}(\operatorname{\mathcal{M}})}$ $\leq\Big\|F_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}|f(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}F_{\epsilon}^{-\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\Big\|G_{\epsilon}^{-\frac{1}{2}}\Big(\int_{\Sigma}|g(\omega)|^{2}{\rm d}\omega\Big)^{\frac{1}{2}}G_{\epsilon}^{-\frac{1}{2}}\Big\|_{L_{\infty}(\operatorname{\mathcal{M}})}$
$=\big\|F_{\epsilon}^{-\frac{1}{2}}F_{0}F_{\epsilon}^{-\frac{1}{2}}\big\|_{L_{\infty}(\operatorname{\mathcal{M}})}\big\|G_{\epsilon}^{-\frac{1}{2}}G_{0}G_{\epsilon}^{-\frac{1}{2}}\big\|_{L_{\infty}(\operatorname{\mathcal{M}})}$
$\leq 1,$
[…]
$\int_{\Sigma}f^{*}(\omega)g(\omega){\rm d}\omega=F_{0}^{\frac{1}{2}}mG_{0}^{\frac{1}{2}},$
which is the desired factorization. ∎
We refer the reader to [19] for more information about Hilbert valued operator spaces.
Acknowledgements
We thank L. Cadilhac for his comment about the $L_{q}$-mean Dini condition, which helped us improve this paper.